### v1

# Factorisation conditions and causality for local measurements in QFT

Abstract
Quantum operations that are perfectly admissible in non-relativistic quantum theory can enable signalling between spacelike separated regions when naively imported into quantum field theory (QFT). Prominent examples of such “impossible measurements”, in the sense of Sorkin, include certain unitary kicks and projective measurements. It is generally accepted that only those quantum operations whose physical implementation arises from a fully relativistically covariant interaction, between the quantum field and a suitable probe, should be regarded as admissible. While this idea has been realised at the level of abstract algebraic QFT, or via particular measurement models, there is still no general set of operational criteria characterising which measurements are physically implementable. In this work we adopt the local $S$-matrix formalism, and make use of a hierarchy of factorisation conditions that exclude both superluminal signalling and retrocausality, thereby providing such a criterion. Realising the local $S$-matrices through explicit interactions between smeared field operators and a pointer degree of freedom, we further derive local causality conditions for the induced Kraus operators, which guarantee the absence of signalling in “impossible measurement” scenarios. Finally, we show that the accuracy with which local field observables can be measured is fundamentally limited by the retarded propagator of the field, which also plays an essential role in a factorisation identity we prove for the field Kraus operators.
I Introduction
The topic of impossible measurements [sorkin1993impossible] has received renewed attention over the past few years. A multitude of approaches have been developed to address the impossible measurement scenarios (see [papageorgiou2024eliminating] and references therein). The clear moral that has emerged is that the set of quantum mechanical operations, or channels, typically allowed in quantum information is far too large when brought to bear on relativistic theories. In the other direction, when one formulates local measurement theory for relativistic quantum field theory (QFT), the resulting channels do not necessarily include the most familiar ones from quantum information.
A straightforward implication is that the quantum informational machinery of states and operations cannot be straightforwardly imported into QFT. In fact, the established empirical success of QFT relies on the asymptotic scattering paradigm, which is not suitable for describing local operations, i.e. operations over bounded regions of spacetime (for historical context see [Fraser2023]). Sorkin’s examples of impossible operations [sorkin1993impossible], and recent refinements [Borsten:2019cpc], showed that operations built out of local operators (i.e. local Kraus operators) generally allow superluminal signalling. Specifically, this includes ideal measurements of many important observables [Benincasa_2014, albertini2023ideal, PhysRevD.105.025003], leading to a number of papers suggesting non-ideal measurement channels, both abstract [PhysRevD.105.025003, Gisin2024towardsmeasurement, albertini2023ideal, oeckl2024spectraldecompositionfieldoperators] and concrete i.e. derived from some purified model [mandrysch2024quantumfieldmeasurementsfewsterverch]. The former use a number of no-signalling conditions on the abstract channels or their Kraus operators, while the latter rely on properties of the purified dynamics.
From the opposite direction, a rigorous framework for quantum measurement theory in algebraic QFT has been proposed by Fewster and Verch (FV framework)[fewster2020quantum, PhysRevD.103.025017, Fewster2023], based on purified dynamics. Together, these results point towards a general theory of measurement and operations in QFT. However, there is still no general set of operational criteria, in terms of e.g. Kraus operators and or their purifications, that characterises which measurement maps and operations are physically implementable. A fundamental step towards this would be a “causal Stinespring theorem”, an idea on which we comment at the end of this paper.
The main results of the paper are as follows. We adapt the local $S$-matrix formalism due to Bogoliubov [pittphilsci24053] to the setting of Sorkin’s impossible operations. We demonstrate which of the factorisation conditions are necessary or sufficient for blocking impossible operations, faster than light signalling and retrocausality. In particular, we demonstrate the sufficiency of the strongest factorisation condition, continuous additivity of the local $S$-matrix
$S[f]=S[f_{+}]S[f_{-}]$ (1)
where $f=f_{+}+f_{-}$, where we have split the support of $f$ in past/future ($_{-}/f+f_{+}$ respectively) with respect to any spacelike hypersurface that goes through the support of $f$. This property allows the decomposition of a local operation with respect to any spacelike hypersurface, but it might not be available for certain interactions. For this reason, we also demonstrate the sufficiency of a weaker factorisation condition, known as the Hammerstein property of the local scattering maps.
Further, we examine the consequences of these factorisation conditions for the induced operations on the field’s Hilbert space, after tracing out the probe system, defined by means of Kraus operators $\{\Pi^{\psi}_{q}[f]\}$ where $\psi,q$ denote states in the probe’s Hilbert space. We show that continuous additivity when combined with the assumption that the $S$-matrix implements quantum control, leads to a no-signalling condition on the associated Kraus operators,
$\Pi_{q}^{\psi}[f]=\int\mathrm{d}s\,\Pi_{q-s}^{\psi_{+}}[f_{+}]\Pi_{s}^{\psi_{-}}[f_{-}].$ (2)
Finally, we use an exactly solvable concrete measurement model that can seen as a relativistic von Neumann measurement model, where a local pointer variable is linearly coupled to a smeared scalar field operator, as $\phi(f)\otimes P$ [PhysRevD.109.065024]. We find that the $S$-matrices and Kraus operators obey the conditions (1) and (2), and so this forms a concrete example of the link between the two representations of channels and no-signalling conditions. Further, by solving the local $S$-matrix model we obtain an explicit class of measurement channels on the field whose sharpness is fundamentally limited by the retarded field propagation within the support of $f$, thereby imposing an intrinsic bound on the accuracy of local field measurements. These channels coincide with the abstract continuous measurement channels studied in [PhysRevD.105.025003, albertini2023ideal, oeckl2024spectraldecompositionfieldoperators], once this new sharpness constraint is taken into account.
The paper is structured as follows: In section II we briefly introduce standard measurement theory and a minimal (algebraic) QFT set-up. We review an abstract notion of operation in the QFT set-up, and briefly present the challenges posed by the impossible measurements scenario, using a simple example of impossible unitary operations introduced in [PhysRevD.105.025003]. In section III we introduce the local $S$-matrix formalism, to discuss the properties of relativistically valid unitary operations in QFT that can (in principle) represent local (not asymptotic) interactions. We take the local $S$-matrix to represent the coupling of the field to a probe system [hellwigkraus1970b] and we examine the operational consequences of a hierarchy of factorisation properties of the $S$-matrix. In section IV we explicitly introduce field operations induced by probes, and the induced Kraus operators in the field’s Hilbert space that are fixed by the local $S$-matrix and the characteristics of the probe system. We consider a simple model for measurements of $\phi(f)$, and consider the limitations to the accuracy of field measurements that come from the causal (retarded) field propagation within the support of $f$. We show how this model relates to the families of Kraus operators for field measurements introduced in Albertini and Jubb [albertini2023ideal] and Oeckl [oeckl2024spectraldecompositionfieldoperators]. Specifically, the result that the sharpness of a measurement of $\phi(f)$ is bounded from below by $\Delta_{\textsc{r}}(f,f)^{1/2}$, i.e. the square-root of the retarded Green’s function smeared against $f$, does not appear in previously suggested measurement maps. In section V we derive no-signalling properties of the induced Kraus operators that can ensure no superluminal signalling and no retrocausality. We discuss the necessity/sufficiency of these conditions and for what type of models they are expected to hold. Finally, we discuss the implications of this model for measurements of more general QFT observables and how this analysis provides hints for relativistic dilation theorems in QFT.
II Local operations in QFT
Perhaps surprisingly, local measurement theory for QFT is still being developed today. This is because the standard predictions and experiments of QFT take the form of asymptotic scattering amplitudes, yet this is not sufficient for modelling local experiments in a relativistic setting. The main obstacle in extracting local probabilistic predictions from QFT theory is that standard quantum mechanical prescriptions for ideal, or projective, measurements and unitary kicks cannot be straightforwardly applied. These operations, when extended to QFT, can lead to signalling across spacelike separated regions. The problem of impossible operations, which was first introduced by Sorkin [sorkin1993impossible], was meant to highlight the friction between no-signalling and the quantum operations that are typically assumed in quantum mechanics. For examples see [Benincasa_2014] and more recently [PhysRevD.105.025003, albertini2023ideal].
In this section we briefly review the basics of standard quantum measurement theory, we introduce a minimal (algebraic) framework for QFT and we discuss the problem of ‘impossible measurements’. We also briefly review the Kraus and Stinespring representations of channels, which will appear in our analysis of QFT channels.
II.1 Standard measurement theory
Quantum mechanics has a standard measurement theory, see [busch_book_2016, busch_standard_1996]. Despite ongoing ambiguity regarding what precisely qualifies certain interactions as measurements – a central aspect of the quantum measurement problem – there exist standard prescriptions for extracting probabilistic predictions for any measurement or general operation defined within the theory. In this subsection we recall the basic ideas of measurement theory, in preparation for application to the QFT setup.
A projective measurement of an observable $A$, represented as a self-adjoint operator in a Hilbert space $\mathcal{H}$, with spectral decomposition
$A=\sum_{j}a_{j}P_{j},$ (3)
involves a set of orthogonal projectors $P_{j}$ corresponding to the possible outcomes $a_{j}$. If the system is initially in a state $\rho$, and the outcome $a_{j}$ is obtained, the post-measurement state is given by the Lüders rule in its selective form:
$\rho\mapsto\frac{P_{j}\rho P_{j}}{\mathrm{Tr}(P_{j}\rho)},$ (4)
where the denominator is the probability of that outcome occurring. If, however, the measurement is performed without recording which outcome occurred—a non-selective measurement—the state updates according to the trace-preserving map
$\rho\mapsto\sum_{j}P_{j}\rho P_{j},$ (5)
which we call an ideal non-selective update. A projective measurement is called ideal, or repeatable.
An observable defines a family of projection operators, called a projection valued measure (PVM), through its spectral decomposition. For an ideal measurement, the elements of the PVM directly relate the probabilities of an outcome, by
$p_{i}(\rho)=\text{Tr}(P_{i}\rho).$ (6)
Here we can see the repeatability. The probability of getting $i$ a second time, given the outcome of the first measurement was $i$ is given by
$p_{i|i}(\rho):=$ $p_{i}(\rho_{i})=\text{Tr}(P_{i}\rho_{i})$
$=$ $\frac{\text{Tr}(P_{i}P_{i}\rho P_{i})}{\mathrm{Tr}(P_{i}\rho)}=\frac{\text{Tr}(P_{i}^{3}\rho)}{\mathrm{Tr}(P_{i}\rho)}=1,$ (7)
where we have used the cyclicity of the trace to move all projections $P_{i}$ together, and then used that $P_{i}^{2}=P_{i}$.
A non-ideal, i.e. non repeatable, measurement must be described by a set of positive operators $E_{i}$ that do not form a complete orthogonal set of projections, but do still compute the probabilities of different outcomes,
$p_{i}=\text{Tr}(E_{i}\rho).$ (8)
Such a family, which sums to the identity, is called a positive operator valued measure (POVM).
The non-selective update (5) corresponds to a completely positive trace-preserving (CPTP) map acting on the system. More generally, any such CPTP map can be represented as arising from a unitary interaction between the system and an ancillary probe (or environment), followed by tracing out the probe. That is, for every CPTP operation $\tilde{\Phi}$ on the system, there exists a unitary $U$ acting on the joint system–probe space and an initial probe state $\sigma$ such that
$\tilde{\Phi}(\rho)=\mathrm{Tr}_{\text{probe}}\big[U(\rho\otimes\sigma)U^{\dagger}\big],$ (9)
[…]
$\rho_{i}=\frac{\Pi_{i}\rho\Pi^{\dagger}_{i}}{p_{i}},$ (10)
where $p_{i}$ is given by (8). Summing over all possible probe outcomes, the non-selective update map becomes
$\tilde{\Phi}(\rho)=\sum_{i}p_{i}\rho_{i}=\sum_{i}\Pi_{i}\rho\Pi_{i}^{\dagger},\;$ (11)
which is equivalent to (9) by the Kraus representation theorem.
We can also define the channel in the Heisenberg picture, acting on $\mathcal{B}(\mathcal{H})$, as
$\Phi(O)=\sum_{i}\Pi_{i}^{\dagger}O\Pi_{i}.$ (12)
$\Phi$ is a unital completely positive channel (UCP), and is dual to $\tilde{\Phi}$ in the sense
$\text{tr}(\rho\Phi(O))=\text{tr}(\tilde{\Phi}(\rho)O),$ (13)
[…]
In QFT it is natural to use the (dual) Heisenberg picture of the channels. Note that in this paper we will mostly consider Kraus-representable non-selective, i.e., UCP channels on the field.
II.2 Set-up
Below we introduce the set-up and we review an example from [PhysRevD.105.025003] that is most relevant for the measurement model that we will consider in section IV.2. We focus on the real scalar field. Finally, we introduce the notion of a local operation, which is key for our understanding of Sorkin’s impossible operations.
Consider a canonically quantised real scalar field theory on Minkowski spacetime $\mathcal{M}$. The field operator $\phi(\mathsf{x})$ at a point $\mathsf{x}\in\mathcal{M}$ is an operator valued distribution and so in order to form an operator algebra we define smeared field operators
$\phi(f)=\int_{\mathcal{M}}\text{dVol }f(\mathsf{x})\phi(\mathsf{x}),$ (15)
for any $f\in C^{\infty}_{0}(\mathcal{M})$, which is the set of all smooth, compactly supported (complex-valued) test functions on $\mathcal{M}$. We further define
$\phi(f),\phi(g)]=i\int_{\mathcal{M}^{2}}\text{d}^{2}\text{Vol }\Delta(\mathsf{x},\mathsf{y})f(\mathsf{x})g(\mathsf{y}):=i\Delta(f,g),$ (16)
where $\Delta(\mathsf{x},\mathsf{y})$ is the usual Pauli-Jordan form
$\Delta(\mathsf{x},\mathsf{y})=\Delta_{\textsc{a}}(\mathsf{x},\mathsf{y})-\Delta_{\textsc{r}}(\mathsf{x},\mathsf{y})$ (17)
where $\Delta_{\textsc{a},\textsc{r}}$ the advanced and the retarded Green’s functions respectively. Note that $\Delta(f,g)=0$ if $\operatorname{supp}f,\operatorname{supp}g$ are spacelike. The commutation relation Eq (16) can be viewed as a condition that must be satisfied by the generators of the local algebras.
In this real scalar field model, abstract local algebras $\mathcal{A}(\text{R})$ can be formed by finite degree polynomials of $1$ and $\phi(f)$ for $\operatorname{supp}{f}\subset\text{R}$. When working in Minkowski spacetime, we can pick the vacuum state $\omega:\mathcal{A}(\mathcal{M})\rightarrow\mathbb{C}$ and construct the Gelfand–Naimark–Segal (GNS) representation $\mathcal{H}$. From now on we will consider all operators to be acting on this space.
The Weyl operators $W(f)=e^{i\phi(f)}$ form a representation of the canonical commutation algebra,
$W(f)W(g)=W(f+g)e^{-i\Delta(f,g)/2}.$ (18)
Taking the Weyl operators with $\operatorname{supp}{f}\subset$ R as generators we can construct a Weyl algebra for each region R, denoted $\mathcal{W}(\text{R})$. Finally, we can take the double commutant $\mathfrak{A}(\text{R}):=\mathcal{W}(\text{R})^{\prime\prime}$—where the commutant of $S\subseteq\mathcal{B}(\mathcal{H})$ is the set of all operators which commute with all elements of SS–––which associates a concrete von Neumann algebra to every region R. We consider an operator $A$ to be localised [^1] AA is localised to R if $A$ is affiliated to $\mathfrak{A}(\text{R})$. to R if $A$ commutes with all operators spacelike to R, i.e. $A,\mathfrak{A}(\text{R}^{\prime})]=0$ where $\text{R}^{\prime}$ is the set of all points spacelike to R. An example is $\phi(f)$, which is local to $\operatorname{supp}{f}\subset\text{R}$, despite being unbounded and thus not in $\mathfrak{A}(\text{R})$. It follows that any bounded function of $\phi(f)$, defined by the spectral calculus, is in $\mathfrak{A}(\text{R})$. For more details on this construction, see [fewster2019algebraic]. Using the properties of the commutator, it follows, from strong commutativity of the fields [Borchers:1992rq], that spacelike algebras commute,
$\mathfrak{A}(\text{R}),\mathfrak{A}(\text{R}^{\prime})]=0,$ (19)
where regions $R$ and $R^{\prime}$ are space-like separated. More generally, in axiomatic approaches to QFT (e.g. algebraic QFT [haag_kastler_1964]) one postulates the axiom of microcausality, namely that spacelike separated algebras of observables commute. This is a kinematical ‘separability’ assumption that, roughly, splits the field degrees of freedom between causally disconnected parts [earman2014relativistic]. Note that microcausality alone does not have an a priori operational consequence [Ruetsche2021, Calderon2022, soulas2025proof] e.g. in terms of signalling. Once a notion of operation is introduced, microcausality alone is generally not sufficient for blocking superluminal signalling (see also [redei_how_2010]). Sorkin showed that [sorkin1993impossible], despite the microcausality assumption, even the most basic local operations in QFT, such as unitary kicks and ideal measurements – when followed by the Lüders post-measurement state update for non-selective measurements – can lead to superluminal signaling between spacelike separated regions (see also [Borsten:2019cpc, PhysRevD.105.025003, oeckl2024spectraldecompositionfieldoperators]).
[…]
for all regions R spacelike to R. This formalises the intuition that a local operation, that in principle is performable within the region R, should not have an effect on regions spacelike to R. This includes unitary kicks, e.g. $e^{iA}$, or non-selective ideal measurements of $A$. Note that selective operations, which we are excluding here, are not local in the sense of Eq (20)  [^3] Note that Eq (20) is not satisfied for selective operations due to the existence of spacelike correlations in generic QFT states. This is a case that we are not considering, since we are concerned with the signaling properties of trace-preserving operations., but typically this is not considered problematic as long as they cannot enable superluminal signalling.
Below, we will briefly review the issue of impossible operations in QFT, i.e., operations that can enable superluminal signalling. This issue highlights that in QFT there is no obvious link between local observables and local operations, as defined abstractly above.
II.3 Impossible operations
Having introduced a minimal AQFT set-up and defined an abstract notion of local operations, we now focus on formalising and presenting Sorkin’s impossible operations. This problem highlights that a local operation, i.e. an operation over a local spacetime region, need not be non-signalling, something that can be seen more clearly in a Sorkin-type scenario involving three (or more) regions. This appears in contrast to microcausality, which only refers to pairs of regions. In the following we will adopt the notation that regions (singled out by the support of smearing functions $f,g,h$ see figure 1) will be denoted as A,B,C etc., while local observables will be denoted as $A,B,C$ etc, e.g. $A\in\mathfrak{A}(\text{A})$, where $\mathfrak{A}(\text{A})$ denotes local algebra, introduced in the next section.
The set-up of the Sorkin problem is the following: consider three regions $\text{A},\text{B},\text{C}$ (figure 1) singled out by the supports of the smearing functions $h,f,g$ respectively. The regions are such that with respect to suitable partial ordering $\text{A}\prec\text{B}\prec\text{C}$ (since B is partially in the causal future of A and C is partially in the causal future of B and the partial order is not transitive), but A and C are spacelike separated. Two operations $\Phi_{\textsc{{a,b}}}$ are applied over regions A and B. For region A to not signal to region C, the expectation values of observables in C must not depend on the choice $\lambda$ of which operation $\Phi^{\lambda}_{\textsc{a}}$ is implemented in A. Namely, the condition for no impossible operations (which has previously been considered in [Borsten:2019cpc, deramon2021relativistic, PhysRevD.105.025003, oeckl2024spectraldecompositionfieldoperators]) is simply the following: let $\rho$ be the global state of the field and $C$ an observable over region C, then it should hold that
$\text{tr}\left(\Phi_{\textsc{a}}^{\lambda}(\Phi_{\textsc{b}}(C))\rho\right)=\text{tr}\left(\Phi_{\textsc{b}}(C)\rho\right),\mbox{ }\forall\lambda.$ (21)
Hence, the expectation values of $C$ should not depend on the operation $\Phi_{\textsc{a}}^{\lambda}$ that is implemented over the spacelike separated region A. We explicitly introduce a dependence of the operation $\Phi_{\textsc{a}}$ over some parameter $\lambda$, which could be used to signal to C if the no-signalling condition (21) fails (see example below). If one demands this for all states $\rho$ and for all $C$, this condition can be viewed as a condition on $\Phi_{\textsc{b}}$ given $\Phi^{\lambda}_{\textsc{a}}$.
Figure 1: Bob’s lab (white) is partly in the future of Alice’s (red) and partly in the past of Charlie’s (blue), while Alice and Charlie are strictly spacelike. Relativistic causality then requires that no signals can be sent from Alice to Charlie, i.e. there must be a no-signalling condition on the experimental setup. Here we have labelled the labs by the supports of the functions $f,g,h$, but the overall causal structure is common to all Sorkin scenarios.
II.3.1 Example of signalling and non-signalling unitaries
The most basic example of failure of this condition using unitary operators in QFT was given in [PhysRevD.105.025003] and uses the real scalar field construction given above. Letting
$\Phi^{\lambda}_{\textsc{a}}(\cdot)=e^{i\phi(h_{\lambda})}\cdot e^{-i\phi(h_{\lambda})},$
$\Phi_{\textsc{b}}(\cdot)=e^{i\phi(f)^{2}}\cdot e^{-i\phi(f)^{2}},$
$C=\phi(g)$ (22)
with supports as shown in figure 1, application of Baker–Campbell–Hausdorff formula leads to
$\text{tr}\left(\Phi_{\textsc{a}}^{\lambda}(\Phi_{\textsc{b}}(C))\rho\right)=2\Delta(h_{\lambda},f)\Delta(f,g).$ (23)
The dependence on $\lambda$ means that Alice can signal to Charlie by picking a family of $h_{\lambda}$ such that $\Delta(h_{\lambda},f)$ is a non-trivial function of $\lambda$. If we replaced $\Phi_{\textsc{b}}$ by
$\Phi_{\textsc{b}}=(\cdot)=e^{i\phi(f)}\cdot e^{-i\phi(f)}$ (24)
then
$\text{tr}\left(\Phi_{\textsc{a}}^{\lambda}(\Phi_{\textsc{b}}(C))\rho\right)=\Delta(f,g).$ (25)
The failure of the no-signalling condition (21) in Eq (23) can be traced back to the fact that $\Phi_{\textsc{b}}(C)$ will not commute with $A$ in general, even though $[A,C]=0$ by microcausality of operators in AQFT [Borsten:2019cpc], while commutation does hold for Eq (25).
Then, it is natural to suspect that the reason why $\Phi_{\textsc{b}}=e^{i\phi(f)^{2}}\cdot e^{-i\phi(f)^{2}}$ is problematic is that $\phi(f)^{2}$ is a non-local operator (given by a double integration over spacetime in this case). Further, one can speculate that the property
$e^{i\phi(f+f^{\prime})}=e^{i\phi(f)}e^{i\phi(f^{\prime})}e^{i\Delta(f,f^{\prime})/2}$ (26)
i.e. that we can split the unitary up to a phase, something that does not hold for the squared field unitary, can be used to rule out impossible operations. Intuitively, this is because it implies that $\Phi_{\textsc{b}}$ can be split into causally ordered parts, which can be chosen so that the part in the future of A drops out the relevant expectation values. As we shall see in section III this type of factorisation property holds for a large and quite general class of unitaries and indeed rules out impossible operations.
III Local $S$-matrix formalism
In this section we will consider a broad set of valid local unitaries in QFT, and their properties. These local unitaries can be interpreted as local operations, and we find that there is a natural causal condition on them that ensures they define non-signalling channels.
Bogoliubov’s local $S$-matrix  [^4] The first attempt to formulate relativistic quantum theory in terms of the $S$-matrix is due to Heisenberg, who introduced a ‘global’ $S$-matrix satisfying unitarity and Lorentz invariance. The local $S$-matrix approach was later introduced by Stueckelberg and Bogoliubov, reintroducing local degrees of freedom and demanding additional causality conditions [pittphilsci24053]. formalism [bogoliubov1960introduction] is a local and covariant approach to describing interactions in QFT, and so is a natural place for us to start. It plays a major role in perturbative algebraic QFT (see [rejzner2019localitycausalityperturbativeaqft] and references within). A local $S$-matrix will be denoted as $S[f]$ where $f$ a compactly supported smearing function. In some cases, when the adiabatic limit can be taken $f(\mathsf{x})\rightarrow f_{\text{const}}$, we end up with the usual global (asymptotic) $S$-matrix familiar from textbook QFT, where coupling terms are spacetime independent constants e.g. $\mathcal{L}_{I}(\mathsf{x})=-f_{\text{const}}\phi(\mathsf{x})^{4}$, and which map between the in and out Hilbert spaces. While the local $S$-matrix formalism was originally conceived, and still used in perturbative AQFT (pAQFT), to define scattering interactions in cases where the asymptotic $S$-matrix might not be well defined, it also provides a useful model for more general localised interactions, which may include scenarios beyond scattering. Further, as we will see, it comes endowed with a hierarchy of causal conditions which are very well suited for studying causality in finite time processes. For these reasons, we will be concerned only with the local case, where the support of $f$ is bounded.
Figure 2: The region $R_{\text{in}}=\mathcal{M}\backslash J^{+}(R)\;(R_{\text{out}}=\mathcal{M}\backslash J^{-}(R))$ (where $\mathcal{M}$ is the spacetime, and $J^{\pm}(R)$ refer to the red and blue regions respectively) for a region $R$, given by the spacelike white regions and the blue (red) region. The out region is the set of all points $\mathsf{x}$ for which $\mathsf{x}\not\preccurlyeq\text{R}$, and likewise the in region is the set of all points $\mathsf{y}$ such that $\text{R}\not\preccurlyeq\mathsf{y}$.
For now, we have not assumed any form for $S$ (e.g. the Dyson expansion). Instead, we assume that $S[f]$ is unitary, and constructed out of operators local to the support of $f$. This leads us to conclude that
$\left[S[f],O\right]=0,\text{ if }O\in\mathfrak{A}(R)$ (27)
for any region $R$ spacelike to $\operatorname{supp}{f}$. Equivalently, we have
$S[f]^{\dagger}OS[f]=O.$ (28)
namely spacelike separated observables remain unchanged. While it may seem that this is sufficient to ensure no-signalling, this is not the case (see next subsection).
One natural way to impose no-signalling, involves zooming into the support of $f$ and make sure that the operation can be decomposed such that future evolution cannot affect past evolution (retrocausality).  [^5] This condition can be imposed on a Dyson expansion of an $S$-matrix by assuming the unsmeared Lagrangian is microcausal. We call continuous additivity [bogoliubov1960introduction] the following property: given a hypersurface $\Sigma$ that intersects the region $\text{supp}f$ (see figure 3) we consider the decomposition $f=f_{+}+f_{-}$ where $\text{supp}f_{\pm}=\text{supp}f\cap J^{\pm}(\Sigma)$. Then
$S[f_{+}]S[f_{-}]=S[f_{+}+f_{-}].$ (29)
This property allows us to decompose the dynamical evolution in past and future evolution with respect to any given foliation.
Figure 3: Continuous additivity of an $S$-matrix corresponds to selection of a (partial) Cauchy surface $\Sigma$ that bisects the support of $f$. The grey region above (below) $\Sigma$ is $J^{+}(\Sigma),\;(J^{+}(\Sigma))$.
III.1 Properties of valid local $S$-matrices
Overall, the relevant properties of the local scattering maps that we will use below are:
  1. 1.
    Unitarity: $SS^{\dagger}=S^{\dagger}S=1$.
  2. 2.
    Normalisation: $S[0]=1$.
  3. 3.
    Locality: An $S$-matrix $S_{\textsc{r}}$ is local to a region R if
    $S_{\textsc{r}},\mathfrak{A}(\text{R}^{\prime})]=0$ (30)
    for all regions R spacelike to R [^6] When the local algebras are von Neumann and Haag duality holds, this condition implies that $S_{\textsc{r}}$ is affilated to $\mathfrak{A}(\text{R})$. Since it is assumed unitary, this is sufficient to imply that $S_{\textsc{r}}\in\mathfrak{A}(\text{R})$..
  4. 4.
    Continuous additivity: Introduced by Bogoliubov, given a local $S$-matrix that depends functorially on a smearing function $f$ such that $S[f]$ is local to $\operatorname{supp}{f}$ and any Cauchy surface $\Sigma$, we can define $_{+}=f\chi_{J_{+}(\Sigma)}, f=f$$_{-}=f\chi_{J_{-}(\Sigma)}$, and have
    $S[f]=S[f_{+}]S[f_{-}].$ (31)
A few comments are in order regarding the meaning of Eq (31), since this property is central in our analysis of (retro)causality [^7] More abstractly, in local operational frameworks (see [oeckl_local_2019]) one demands a similar compositionality property of null probes (that represent dynamics) that share a boundary. . For our purposes, we demand (31) inspired by the causal perturbation theory program888where the condition is normally introduced in order to help defining the local algebras of a perturbative and causal QFT, see [Ilyin:1978gq, Brunetti_2000, rejzner2019localitycausalityperturbativeaqft] and a recent project for non-perturbative AQFT [Buchholz_2020]. In these works, it is analysed in detail how restrictive this assumption is for realistic interactions, which of course goes beyond our purposes., as well as weaker factorisation conditions (see section VII.1) to analyse their operationally meaningful consequences, in the context of signalling, in a simple measurement scenario involving a probe locally coupled to a field. A final comment about (31) is that it is key that we are cutting $f$ on a Cauchy slice, i.e. a spacelike hypersurface, rather than a lightlike hypersurface such as one that coincides with A or C’s lightcones, as factorising the local $S$-matrix on the lightcone would introduce divergences (which can be seen using microlocal analysis [Brunetti_2000]).
Finally, so far our discussion has been limited to a QFT—whether consisting of multiple fields or a single one—in which the scattering matrix is restricted to act only within a single region (B). To analyse information flow between different regions, each of which may involve the application of a distinct scattering matrix, we now extend the framework to encompass multiple regions. For composing distinct (causally orderable) interactions we additionally need to assume the following:
  1. 5.
    Causal factorisation: Given two disjoint regions A, B with B in the out region  [^9] The out region $R_{\text{out}}$ of some region $R\subset M$ is given by $M\backslash J^{-}(R)$. of A (see Fig 2), the evolution is given by an $S$-matrix $S_{\textsc{a+b}}$ which factorises
    $S_{\textsc{a+b}}=S_{\textsc{b}}S_{\textsc{a}}$ (32)
    where $S_{\textsc{a}}\;(S_{\textsc{b}})$ is the $S$-matrix where just region A(B) has non trivial coupling, and is local to A(B).
Note that if the supports are fully spacelike separated then it also holds that $S_{\textsc{b}}S_{\textsc{a}}=S_{\textsc{a}}S_{\textsc{b}}$.
III.1.1 Eliminating the ‘impossible’
To appreciate the operational consequences of the properties listed above, we need to define a valid notion of operation based on the local $S$-matrix. The first step is to enlarge the Hilbert space by including the Hilbert space of probe systems that are coupled to the field through local $S$-matrices (one probe per interaction region). Consider $n$ probes and $n$ regions $\text{R}_{i}$, each of which are disjoint but not necessary all mutually spacelike, given by quantum systems defined on $\mathcal{H}_{\text{QFT}}\bigotimes_{i=1}^{n}\mathcal{H}_{i}$. Each of these probes can be associated to an interaction term
[…]
which is a coupling between the field and one probe, localised to some region $R_{i}$. Approaches to measurement theory in QFT in which the measurement process is modelled as a local scattering process and operations are induced by probes, include the detector-based measurement theory [Polo_Gomez_2022] and the Fewster-Verch (FV) framework in algebraic QFT [fewster2020quantum]. In the detector-based approach, one works with (a general class) of Hamiltonian densities that model the detector-field coupling e.g. mimicking the light-matter interaction, and the local $S$-matrix can be defined as the time-ordered exponential of these interaction Hamiltonians (typically the predictions of the models rely on perturbation theory, using the Dyson expansion of the $S$-matrix). In the algebraic approach, the formulation of the FV framework relies on the structure of abstract *-algebras. In this setting, the interaction between system and probe field is formulated as a morphism $\Theta$ of the uncoupled theory, and it is not necessarily implemented through the adjoint action of a unitary scattering map $S$, namely as $\Theta(\cdot)=S^{\dagger}(\cdot)S$. Note that the approach we are taking here, is more abstract than the first (since it does not rely on particular interaction Hamiltonians and the Dyson expansion), but more concrete than the latter, since we assume the existence of these local scattering maps that implement the scatteting morphism in a Hilbert space.
Partial or exact resolution to the impossible measurements scenario have been offered respectively in the detector-based approach [benincasa2014quantum, deramon2021relativistic] and in the FV-framework [PhysRevD.103.025017]. In either case, an important ingredient of the resolution was an analogue causal factorisation property (32). An abstract version of causal factorisation is assumed to hold in the FV framework  [^10] The abstract version of the causal factorisation condition is that $\Theta_{\textsc{a+b}}=\Theta_{\textsc{b}}\circ\Theta_{\textsc{a}}$, which has to be proven for concrete models in this framework., while in [deramon2021relativistic] it was argued that that this property holds for a general class of Unruh-DeWitt-type detector models coupled locally to a coupled field over disjoint, causally orderable regions. Finally, all the measurement schemes considered in QFT correspond to so-called pre-measurement processes, involving only the unitary interaction between the field and the probe. Strictly speaking, the measurement is completed only once an outcome has been obtained, which would require, in turn, a measurement of the probe and so on, leading to an infinite regress, see [mandrysch2024quantumfieldmeasurementsfewsterverch]. In this work we ignore this final measurement stage and, following common jargon, refer to the pre-measurement stage itself as the “measurement process”.
We should emphasise that in both approaches causal factorisation is assumed (or proven on a case-by-case basis) but, crucially, it is not sufficient for eliminating impossible measurements. More assumptions about the probe-field interactions in the coupling regions have to be made, i.e. the locality of the detector’s local current in the detector model approach [deramon2021relativistic]  [^11] In [deramon2021relativistic] the authors show that (seemingly local) Hamiltonian densities of the form $h(\mathsf{x})=\Lambda(\mathsf{x})J(\mathsf{x})\otimes\phi(\mathsf{x})$, where $J(\mathsf{x})$ is the current associated to a probe, can give rise to impossible measurements if the probe’s dynamical current is not microcausal (i.e. if $J(\mathsf{x}),J(\mathsf{y})]\neq 0$ for $\mathsf{x},\mathsf{y}$ is not spacelike separated) by using the Dyson expansion of $S=\mathcal{T}\text{exp}i\int h(\mathsf{x})\mathrm{d}\mathsf{x}$. , or the time-slice property of the coupled theory in the FV framework  [^12] The exact relation between continuous additivity and other dynamical axioms e.g. the time-slice property in the formal setting of (p)AQFT goes way beyond our purposes (for related results see e.g. [ruep_thesis_2022]). [PhysRevD.103.025017]. Our claim is that Eq (31) captures the properties of a causal probe-field coupling that arises from the additional assumptions needed in the concrete and abstract frameworks, while making it clear that this can be directly related to internal factorisation properties of the interactions. Specifically, we will see that the ability to appropriately factorise the interaction region into causally ordered subregions is sufficient for ruling out impossible, i.e. signalling, operations. In the next section, we will see that it is not necessary, since a weaker factorisation property can be used.
To see that (31) is sufficient, let us consider a Sorkin scenario, where we assume disjoint operational factorisation between the coupling regions, and continuous additivity for Bob’s, see figure 4. The total $S$-matrix is given by
$S_{\textsc{a+b}}=S_{\textsc{b}}[f]S_{\textsc{a}}=S_{\textsc{b}}[f_{+}]S_{\textsc{b}}[f_{-}]S_{\textsc{a}}$ (34)
and so
$\text{tr}(S_{\textsc{a+b}}^{\dagger}$ $CS_{\textsc{a+b}}\rho)$
$=$ $\text{tr}(S_{\textsc{a}}^{\dagger}S_{\textsc{b}}[f_{-}]^{\dagger}S_{\textsc{b}}[f_{+}]^{\dagger}CS_{\textsc{b}}[f_{+}]S_{\textsc{b}}[f_{-}]S_{\textsc{a}}\rho)$
$=$ $\text{tr}(S_{\textsc{a}}^{\dagger}S_{\textsc{b}}[f_{-}]^{\dagger}CS_{\textsc{b}}[f_{-}]S_{\textsc{a}}\rho)$
$=$ $\text{tr}(S_{\textsc{b}}[f_{-}]^{\dagger}CS_{\textsc{b}}[f_{-}]\rho)$ (35)
where we have used the locality properties—i.e. $S[f]$ is localised to $\operatorname{supp}{f}$ and $S_{R}$ is localised to region R for R==A,B— to commute $S[f_{+}]$ with $C$ and $S_{\textsc{a}}$ with $S[f_{-}]$ and $C$, see again figure4. Since there is no trace of Alice’s influence in the final expression, there can be no superluminal signalling from A to C.
Importantly, also note that only the part of Bob’s coupling that is in contact with Charlie’s causal past, i.e. $S[f_{-}]$ survives in the final expression, while $S[f_{+}]$ which is fully spacelike separated from Charlie has dropped out. Yet, for any fixed splitting of region B, the support of $f_{-}$ contains points that are spacelike separated from C, as the splitting is along a spacelike Cauchy surface. In principle, changes to the $S$-matrix at these points should not affect expectation values at C. This observation will become important when we introduce the notion of effective Kraus operator in section V.
Finally, note that introducing the probes above was only the first step towards defining a notion of operation in the next sections. We will do this by suitably taking partial traces over the probes’ degrees of freedom, to induce an operation on the Hilbert space of the field, while so far we are only describing the full field-probes dynamics and calculating expectation values of field observables.
Figure 4: A Sorkin scenario where the dynamics of the interior region can be split as given in Eq (31). Note that $\Sigma$ is chosen such that $\operatorname{supp}{f_{-}}$ is spacelike to A and $\operatorname{supp}{f_{+}}$ is spacelike to C.
III.2 Hierarchy of factorisation conditions
We have introduced two notions of factorisations of $S$-matrices, causal factorisation and continuous additivity, that respectively correspond to composing disjoint interactions or decomposing a single interaction. Overall, there is a hierarchy of spacetime factorisation properties, as explained in [Brunetti_2000]. In this section, we will briefly outline them here for completeness, and generalise our proof of no impossible operations to a weaker condition.
Firstly, we note that (29) is quite a stringent condition on the local $S$-matrix. This condition might fail for certain singular interactions (in certain spacetime dimensions). For this reason, here we will also consider weaker factorisation conditions that are more widely assumed to hold for local interactions.
One such weaker condition, called the Hammerstein property, is typically assumed in causal perturbation theory literature [rejzner2019localitycausalityperturbativeaqft], where it is taken as continuous additivity Eq (31) may be too strong for very singular interactions [Brunetti_2000]. The Hammerstein property is that
$S[f_{1}+f_{2}+f_{3}]=S[f_{1}+f_{2}]S[f_{2}]^{-1}S[f_{2}+f_{3}]$ (36)
where $f_{1},f_{3}$ are separated by a hypersurface and $f_{2}$ is unrestricted (and it can potentially overlap with $f_{1,3}$). Continuous additivity
$S[f]=S[f_{+}]S[f_{-}]$ (37)
where $f=f_{+}+f_{-}$ and $f_{\pm}$ need not be smooth at the hypersurface $\Sigma$, in turn implies the weaker condition (36). To see this, we split $f$ into $f_{+}+f_{0}+f_{-}$ described for Eq (36), and apply condition Eq (29),
$S[f]=$ $S[f_{+}+f_{0}-f_{0}+f_{0}+f_{-}]$ (38)
$=$ $S[f_{+}+f_{0}]S[f_{0}]^{\dagger}S[f_{0}+f_{-}]$ (39)
where we have used the stronger condition Eq (29) to factorise, and used that $S[0]=S[f]S[f]^{\dagger}\implies S[f]^{\dagger}=S[-f]$.
Importantly, the Hammerstein property is also sufficient to rule out Sorkin problems, as we explicitly show in detail in the appendix VII.1 and sketch here.
Figure 5: The spacetime factorisation of $f$ into $f_{+}$ (red) $f_{0}$ (white) and $f_{-}$ (blue), where we note the disjoint and time ordered supports of $f_{\pm}$, and the support overlaps of $f_{0}$ with $f_{\pm}$. Key to the proof is that the support of $f_{0}$ is spacelike separated from both A and C.
Looking at Figure 5, we can see a similar scenario to the proof for the stronger factorisation condition. As there, the supports of $f_{\pm}$ are spacelike to C,A respectively, and now $f_{0}$ is spacelike to A and C. This implies that $S[f_{+}+f_{0}]S[f_{0}]^{\dagger}$ commutes with $C$ and $S[f_{-}+f_{0}]^{\dagger}CS[f_{-}+f_{0}]$ is localised spacelike to A. Following a very similar derivation to Eq (III.1.1) demonstrates that there is no signalling in a Sorkin scenario where $\Phi_{\textsc{b}}$ is induced by an $S$-matrix satisfying the Hammerstein property  [^13] Resolutions of the Sorkin problem in the abstract algebraic setting of the FV framework rely on the time-slice property of the coupled theory [PhysRevD.103.025017, papageorgiou2024eliminating]. At the level of the local $S$-matrix it has been shown that the time-slice property can be derived only from (36) and normalisation $S(0)=\mathsf{1}$ [chilian2009time], which explains why we were able to demonstrate the link between (36) and no impossible operations in appendix VII.1.. For the full details, see appendix VII.1.
Finally, (36) implies the weakest factorisation condition (for $f_{2}$ being the zero function), that
$S[f_{1}+f_{3}]=S[f_{1}]S[f_{3}]$ (40)
where $f_{1},f_{3}$ are again separated by a Cauchy surface but have disjoint supports, e.g.
$\overline{\operatorname{supp}{f_{1}}}\cap\overline{\operatorname{supp}{f_{2}}}=\emptyset.$ (41)
Note that the weakest factorisation condition, causal factorisation, (40) is necessary but not sufficient for resolving the Sorkin problem, while the strongest (32) completely rules out Sorkin’s impossible operations. So what is the role of the middle condition (36)? This property appears as a major component of causal perturbation theory, where for very singular interactions the strongest condition may not hold. Since the concrete model for field measurements that we consider in the next section satisfies the stronger, and simpler, condition (31), we will utilise that instead for translating it into properties of induced Kraus operators.
IV Operations induced by probes
In quantum mechanics, every unital, i.e. maps the identity to itself, CP map (UCP map) – that is, every non-selective operation on a quantum system – can be understood as arising from a unitary interaction between the system and an ancillary system, or probe, followed by discarding (tracing out) the probe. This is a consequence of the Stinespring dilation theorem, which ensures that any UCP transformation of the system can be “purified” by embedding it into a larger closed system that evolves unitarily. In the same vein, we will consider in this section the realisation of non-selective operations in QFT through interactions between the system (i.e., the field) and a probe. Then, a way to argue that unitary “kicks” which violate no-signalling are not physically realisable is to show that they cannot be implemented by any valid local interaction between the system and a probe (see also [muchverch]).
Below, we define local Kraus operators as induced by a suitable local coupling $S[f]$ between the field and the probe, and we provide an example of a concrete model for field measurements. In the following section, we derive no-signalling conditions for the associated Kraus operators.
IV.1 Induced Kraus operators in QFT
Kraus operators provide a way to describe non-unitary channels using operators $\Pi_{q}\in\mathcal{B}(\mathcal{H})$. In this section we briefly describe their properties in QFT.
The attempts of modelling local operations in QFT required a departure from the asymptotic scattering paradigm. In a series of papers [hellwigkraus1969, hellwigkraus1970b] Hellwig and Kraus model relativistic quantum measurement as a local scattering process. Their motivation for introducing the so-called Kraus operators, as induced through the dynamical coupling to some probe, was to explicate the operational interpretation of the local algebras that was implicit in the axiomatic approaches to QFT [haag_kastler_1964]. Their measurement scheme can be viewed as predecessor of recently proposed measurement frameworks in QFT [Polo_Gomez_2022, fewster2020quantum] that involve some (abstract or concrete) representation of dynamics between system and probe (for historical context see [Fraser2023]).
Consider the combined Hilbert space $\mathcal{H}^{(s)}\otimes\mathcal{H}^{(p)}$ of a system (field) with Hilbert space $\mathcal{H}^{(s)}$ and some probe with Hilbert space $\mathcal{H}^{(p)}$. Using interaction picture the final result of the probe-system interaction can be described through a unitary operator $S$ on the product space $\mathcal{H}^{(s)}\otimes\mathcal{H}^{(p)}$. We will assume that the coupling $S$ is supported over a bounded spacetime region, i.e., $S=S[f]$ where $f=f(\mathsf{x})$ is a spacetime smearing function that is compactly supported.
We will only use the the dual update map, acting on the operators rather than the states. This is most natural as the QFT operators we are working with are in the Heisenberg picture. The update rule is now given by
$O\mapsto\int\mathrm{d}q{\Pi^{\psi}_{q}[f]}^{\dagger}O\Pi^{\psi}_{q}[f]=:\Phi[f](O)$ (42)
where we have included the functional dependence on the spacetime smearing function $f$, and the Kraus operators for the field associated to $S[f]$ are given by
$\Pi^{\psi}_{q}[f]:=\braket{q|S[f]|\psi},$ (43)
[…]
Kraus operators in scattering-based frameworks for local measurement in QFT have been derived previously in [SmithAlexander2017, Polo_Gomez_2022, PhysRevD.109.065024] for concrete Unruh-DeWitt-type models, and more recently in [mandrysch2024quantumfieldmeasurementsfewsterverch] for a concrete representation of the FV framework in a Hilbert space. Nevertheless, the operational consequences of the factorisation conditions at the level of Kraus operators have not been investigated. Here we investigate how these properties are captured by valid local scattering matrices and their (de)composition properties. Further, for certain concrete model for measurement of field amplitudes, we link these properties to a no-signalling property for the resulting Kraus operators.
IV.2 A concrete model for local field measurements
We will use a simple model for field measurements (previously considered in [PhysRevD.109.065024]) where a scalar quantum field $\phi$ in Minkowski spacetime is coupled to a free pointer via $P$, with $[Q,P]=i$. We specify the following interaction
$\mathcal{L}_{I}[f]=f(\mathsf{x})\phi(\mathsf{x})\otimes P.$ (44)
Since we can pick any free Hamiltonian for the pointer that commutes with the interaction without much computational difficulty, we could in principle pick that of a relativistic scalar particle. However, to avoid any conceptual problems arising from interpreting the pointer as a relativistic particle—given the well known problems of causality and localisation in relativistic QM (e.g. Hegerfeldt’s [hegerfeldt2001particle] and Malament’s [malament1996defense] theorems)—we elect simply to set $H_{\text{pointer}}=0$, viewing the pointer as an abstract, degree of freedom, not tied to spacetime. This has no effect on the field Kraus operators, and the change to the $S$-matrix can be found easily. To emphasise this conceptual choice, we will continue to denote spacetime points by $\mathsf{x}$ etc, while the internal values of the pointer as $q$. The interaction in (44) leads to the following scattering matrix
$S[f]=\mathcal{T}e^{-i\phi(f)\otimes P}$ (45)
which we can solve exactly via the Magnus expansion [PhysRevD.109.065024] to give
$|S[f]=\int\mathrm{d}p\,e^{-ip\phi(f)}e^{-ip^{2}\Delta_{\textsc{r}}(f,f)/2}\otimes\ket{p}\bra{p}$ (46)
where $\Delta_{\textsc{r}}$ is the retarded propagator (see Appendix VII.2 for a derivation) and
$\Delta_{\textsc{r}}(f,f)=\int\mathrm{d}\mathsf{x}\mathrm{d}\mathsf{y}f(\mathsf{x})\Delta_{\textsc{r}}(\mathsf{x},\mathsf{y})f(\mathsf{y}).$ (47)
After picking a probe state $\psi(q)$ on the basis of the pointer variable $Q$, using Eq (43) we can compute the Kraus operators for the field [PhysRevD.109.065024]
$\Pi_{q}^{\psi}[f]=$ $\frac{1}{\sqrt{2\pi}}\int\mathrm{d}p\,\tilde{\psi}(p)$
$\times\text{exp}\left(-\frac{i}{2}p^{2}\Delta_{\textsc{r}}(f,f)+ip(q-\phi(f))\right),$ (48)
where $\tilde{\psi}(p)=\langle p|\psi\rangle$ is the Fourier transform of $\psi(q)$. Note that the expression (IV.2) has the form of a Fourier transform with an additional phase that depends on $\Delta_{\textsc{r}}(f,f)$, which arises due to the time-ordering operator in Eq (45) (using the Magnus expansion, see Appendix VII.2).  [^14] The retarded propagator is due to the time-ordering operator, using the Magnus expansion for the time-ordered exponential. The deeper reason why we end up only with retarded (and not the advanced) propagation is the implicit time-asymmetry introduced by assuming initial (and not final) separable state. It is precisely the causal retarded propagation within the support of $f$ that this model captures at the level of the Kraus operators, which will play an essential role in analysing the resolution of field measurements, and the no-signalling properties of the Kraus operators in the next section.
As an example, by picking $\psi(q)=G_{\sigma}^{1/2}(q)$ to be the square root of the Gaussian centred at 0 with width $\sigma$, we have the following POVMs
${\Pi^{\sigma}_{q}[f]}^{\dagger}\Pi_{q}^{\sigma}[f]$ $=\frac{1}{\sqrt{2\pi}\tilde{\sigma}}\text{exp}\left[-\frac{\left(q-\phi(f)\right)^{2}}{2\tilde{\sigma}^{2}}\right]$
$=G_{\tilde{\sigma}}(q-\phi(f))$ (49)
where
$\tilde{\sigma}=\sqrt{\frac{\sigma^{2}}{2}+\frac{\Delta_{\textsc{r}}(f,f)^{2}}{2\sigma^{2}}}\geq\Delta_{\textsc{r}}(f,f)^{\frac{1}{2}}$ (50)
Hence, $\tilde{\sigma}$ takes its minimal value of $\Delta_{\textsc{r}}(f,f)^{\frac{1}{2}}$ when $\sigma^{2}=\Delta_{\textsc{r}}(f,f)$. Note that this POVM governs the statistics of the conjugate pointer variable $Q$ after the coupling is ‘off’ [^15] By this, we mean when we are in the out region of $\operatorname{supp}{f}$..
The uncertainty $\tilde{\sigma}$ represents how far the POVM is away from being a projection valued measure (PVM), or in quantum measurement language, how far from sharp the measurement is. It is not hard to see that if we could take $\tilde{\sigma}\rightarrow 0$, then the POVM would become
$\Pi^{\sigma}_{q}[f]^{\dagger}\Pi_{q}^{\sigma}[f]\rightarrow\delta(q-\phi(f))$
$=$ $\int\mathrm{d}\lambda\delta(q-\lambda)dP(\lambda)=P(q)$ (51)
where $P(q)$ is the PVM associated to the possible values of $\phi(f)$ by the spectral theorem.
In [ozawa1993canonical], Ozawa studies the process of taking such a limit, and shows that a measurement is only repeatable in that limit, in the sense that the outcomes of repeated measurements agree. Operationally, they consider allowing the interaction to continue for a longer time, say a factor of $T$, and rescaling the pointer units by $1/T$. It is shown in the case considered in that paper the standard deviation is scaled by $1/T$. This is mathematically the same as scaling the probe $x$ state’s argument by $T$, thus scaling $\sigma$ by $1/T$. We cannot do quite the same here, as we now discuss.
Suppose we take $\sigma$, the probe’s standard deviation in $q$, to zero, then the first term in Eq (50) will indeed vanish, however, the second term will explode, meaning the uncertainty diverges, $\tilde{\sigma}\rightarrow\infty$. The appearance of the propagator term in Eq (50) is a consequence of solving the Dyson series for the $S$-matrix exactly, rather than assuming the interaction Hamiltonian dominates the field’s free Hamiltonian. We can also see that the square root of the propagator bounds the uncertainty from below, and so we must simultaneously reduce $\Delta_{\textsc{r}}(f,f)$ and take the probe state to be sharper. If done such that the inequality in Eq (50) is saturated at each step, then the POVM will become sharper and sharper. However, this requires $\Delta_{\textsc{r}}(f,f)$ be taken smaller. In the next section we interpret this bound further.
IV.2.1 Regions and sharpness
To gain some intuition for the bound (50) on how accurately the field values can be probed locally, we can appeal to the results of [PhysRevD.109.065024], where it is shown that for $f$ given by a Gaussian with spacial width $\sigma_{S}$ and time width $\sigma_{T}$,
$\Delta_{\textsc{r}}(f,f)\sim\begin{cases}\frac{1}{\sigma_{T}\sigma_{S}^{3}}&\text{ when }m\sigma_{T}\gg 1\\ \frac{\sigma_{T}}{\sigma_{S}(\sigma^{2}_{S}+\sigma_{T}^{2})}&\text{ when }m=0.\end{cases}$ (52)
The intuition is that the effect of interacting with the field will disturb it, and those disturbances will propagate following the dispersion relations. When $m\gg\frac{1}{\sigma_{T}}$, those disturbances will travel non-relativistically, and so will not probe the spacial extent of the region $f$, while when $m\sim 0$, they will travel close to or at the speed of light, see figure 6.
The retarded propagator encapsulates how much a perturbation at $\mathsf{x}$ affects the field at $\mathsf{y}$. For all masses, it vanishes if the points are spacelike separated, and if the mass is zero then it diverges when there is a lightlike curve from $\mathsf{x}$ to $\mathsf{y}$, meaning that the perturbation caused by the measurement interaction at $\mathsf{x}$ will propagate along the lightcone and affect the measurement at $\mathsf{y}$. If the region is spacially wide but temporally narrow, i.e. the region is not well causally connected, then not much of the support of $f$ is contained in the future lightcone of any point $\mathsf{x}$, and so there is minimal disturbance. If the smearing is purely spacial (instantaneous interaction) then $\Delta_{\textsc{r}}(f,f)=0$.
Conversely, if the region is spacially narrow and temporally long, then this is flipped on its head, and the disturbance is large, see figure 6. Finally, if the mass of the particle is non-zero, then the retarded propagator is no longer concentrated on the lightcone, as represented by the timelike worldlines in figure 6. In this case, the causal connectivity of the region is less important, and the disturbance instead depends on the overall spacetime volume of the region.
Figure 6: (Left) a region that has a large time extension will contain overlapping futures of points that are at large spacial separations e.g. $J^{+}(P)\cap J^{+}(Q)$, indicated by the darkest grey region. Conversely, if the degrees of freedom of the theory travel much slower than $c$, these overlap regions will not be important, indicated by the two black timelike curves. (Right) A region with small time extension will only have points that are close together have overlapping futures contained in itself: $J^{+}(P)\cap J^{+}(Q)=\emptyset$.
In the $m\gg\frac{1}{\sigma_{T}}$ case, increasing either size of $\sigma_{T},\sigma_{S}$ makes the measurement sharper, while for $m\sim 0$, when $\sigma_{S}^{2}+\sigma_{T}^{2}$ is held constant, the region must be taken to be much wider spacially than temporally in order to get a sharp measurement. We see that concretely modelling a measurement leads to a channel that would not be easily guessed if avoiding signalling was the only concern. Indeed, channels comparable to ours have been guessed, as we will see in the next section, however they generally lack the bound on the sharpness we have found, which comes from the $\Delta_{R}(f,f)$ phase in the $S$-matrix. This phase arose, not from causality constraints, but from explicitly solving for the evolution of the field and pointer. As of such, it goes beyond Sorkin’s impossible measurements, as the channel is still causal without the $\Delta_{R}(f,f)$ phase. Interestingly, while the appearance of $\Delta_{R}(f,f)$ is not necessary for ruling out signalling of the field channel, it is key in the proof of continuous additivity and thus plays a role in the causality of the dilated channel, see appendix VII.3.
IV.2.2 The model as a dilation
Finally, we comment on the connection to previous related work on families of Kraus operators and POVMs for field measurements. Our analysis provides a concrete dilation of the $L^{2}$-maps that defines families of Kraus operators in [albertini2023ideal], defined as
$K^{f}_{\kappa}(q)=\kappa(q-\phi(f)).$ (53)
By picking the internal state of the probe to be
$\psi(q):=\Psi_{f}(\kappa)[q]:=\mathcal{F}^{-1}\left(\mathcal{F}(\kappa)[p]e^{\frac{ip^{2}}{2}\Delta_{\textsc{r}}(f,f)}\right)[q]$ (54)
where $\mathcal{F},\mathcal{F}^{-1}$ are the Fourier and inverse Fourier respectively, we can see that
$K^{f}_{\kappa}(q)=\Pi^{\Psi_{f}(\kappa)}_{q}[f].$ (55)
This means that the ‘shape’ of the Kraus operators is specified by the initial state preparation $\psi$ of the probe. Similarly, there is a connection to the Kraus operators that are defined in [oeckl2024spectraldecompositionfieldoperators] in the case of instantaneous field measurements, that we will comment on in the next section. Lastly, in [mandrysch2024quantumfieldmeasurementsfewsterverch] the authors analyse a linear coupling between two equal mass real scalar field theories, building on [fewster2020quantum, Fewster2023]. By picking a squeezed quasi-local state for the probe field, they show that a Gaussian measurement of the probe leads to an instrument that in the weak coupling limit can be decomposed into a Gaussian measurement and a dephasing channel with certain strength. A full comparison between these results would require letting the masses of both fields vary independently. his is because our pointer is a single abstract degree of freedom, and has trivial internal dynamics, which would appear in a well controlled large mass limit of the probe represented as a quantum field, likely requiring us to reintroduce a probe Hamiltonian. We leave further exploration for future works.
V Causality conditions on the Kraus operators
In this section we will examine the relevance of the continuous additivity Eq (31) for the concrete model of field measurements that we introduced in the previous section. We will demonstrate the model’s consistency with the framework laid out in section III, and derive the no-signalling and compositional properties of its Kraus operators, generalising those found in [oeckl2024spectraldecompositionfieldoperators]. These no-signalling identities ensure that this model does not induce impossible operations on the field.
The fact that an imprint of the causal (retarded) dynamics of the field appears in the expressions of the local scattering map of this model Eq (46) and the corresponding Kraus operators Eq (IV.2) leads us to the following observation: for any Cauchy slice through $f$ (see figure 4) continuous additivity holds for this model, i.e.,
$S[f]=S[f_{+}]S[f_{-}],$ (56)
[…]
Crucially, this has the following consequence for the Kraus operators, which we prove in appendix VII.5- that they obey
$\Pi_{q}^{\psi}[f]=\int\mathrm{d}s\,\Pi_{q-s}^{\psi_{+}}[f_{+}]\Pi_{s}^{\psi_{-}}[f_{-}],$ (57)
where $f=f_{+}\!+f_{-}$ and $\psi=\psi_{+}\star\psi_{-}$, where $\star$ is convolution. This is a way of decomposing a single time-extended measurement process, for different choices of spacelike hypersurfaces going through the region. Note that $\psi^{\pm}$ are not necessarily square normalised and do not represent valid states for the probe, so $\Pi_{q}^{\psi_{+}}[f_{+}],\Pi_{q}^{\psi_{-}}[f_{-}]$ do not (yet) define valid (normalised) Kraus families (see (V.1) in the next subsection).
Identity (57) will be key in showing that the such an operation cannot enable signaling between spacelike separated regions in a Sorkin-type scenario. In appendix VII.5 we prove it for a simple model for $\phi(f)$ measurements, but more generally, it follows from the continuous additivity of $S[f]$ and the fact that $S[f]$ can be seen as a quantum controlled unitary (see subsection V.3). This provides us with a concrete example of the relationship between no-signalling conditions for dilated and undilated maps.
Identity (57) was proved in [oeckl2024spectraldecompositionfieldoperators] for an $\epsilon$-family of POVMs, including the spectral decomposition of $\phi(f)$. in the case of instantaneous operations defined over a Cauchy surface ($f$ being a purely spacial smearing) and it was proved to be sufficient for enforcing no-signalling condition Eq (21) in this case. Note that our results reduce to those in [oeckl2024spectraldecompositionfieldoperators] when we take $\psi$ to be Gaussian with width $\varepsilon$ and $f$ is supported on a spacial hypersurface so that $\Delta_{\textsc{r}}(f,f)=0$.
In our case, our concrete model describes time-extended operations, so (57) is not going to be sufficient to prove no-signalling and we will need the following additional identity [^16] The connection between this work and a generalised, i.e. ‘time extended’, version of Oeckl’s results in [oeckl2024spectraldecompositionfieldoperators] using a covariant path-integral description [oeckl2025causalmeasurementquantumfield] is work in progress, see also [oeckl2025localcompositionalmeasurementsquantum]..
[…]
where
$F_{\psi}(s-s^{\prime})=\int\mathrm{d}p\,|\widetilde{\psi}(p)|^{2}e^{ip(s-s^{\prime})},$ (59)
which we derive in appendix VII.5.
V.1 No impossible operations
To see this, consider a Sorkin scenario where Alice applies $\Phi_{\textsc{a}}$, Bob applies $\Phi_{\textsc{b}}[f]$—where we have been explicit about only Bob’s map’s functional dependence—and Charlie measures the expectation value of $C$. Picking a Cauchy hypersurface $\Sigma$ such that A is in the future and C is in the past (see figure 3), we follow the notation from before for splitting $f=f_{+}+f_{-}$. We then have the following implication
$\text{tr}\left(\Phi_{\textsc{a}}\circ\Phi_{\textsc{b}}[f](C)\rho\right)=$ $\text{tr}\left(\Phi_{\textsc{a}}\circ\int dx\Pi^{\psi}_{q}[f]^{\dagger}C\Pi^{\psi}_{q}[f]\,\rho\right)$
$=$ $\int\mathrm{d}q\mathrm{d}s\mathrm{d}s^{\prime}\text{tr}\left(\Phi_{\textsc{a}}\circ\Pi^{\psi_{-}}_{s}[f_{-}]^{\dagger}\Pi^{\psi_{+}}_{q-s}[f_{+}]^{\dagger}C\Pi^{\psi_{+}}_{q-s^{\prime}}[f_{+}]\Pi^{\psi_{-}}_{s^{\prime}}[f_{-}]\rho\right)$
$=$ $\int\mathrm{d}q\mathrm{d}s\mathrm{d}s^{\prime}\text{tr}\left(C\Pi^{\psi_{-}}_{s^{\prime}}[f_{-}]\rho\Pi^{\psi_{-}}_{s}[f_{-}]^{\dagger}\Phi_{\textsc{a}}\left(\Pi^{\psi_{+}}_{q-s}[f_{+}]^{\dagger}\Pi^{\psi_{+}}_{q-s^{\prime}}[f_{+}]\right)\right)$
$=$ $\int\mathrm{d}s\mathrm{d}s^{\prime}\text{tr}\left(F_{\psi_{+}}(s-s^{\prime})\Pi^{\psi_{-}}_{s}[f_{-}]^{\dagger}C\Pi^{\psi_{-}}_{s^{\prime}}[f_{-}]\rho\right)$ (60)
In the second line above we have used the identity Eq. (57) to split each Kraus operator into those localised before and after the Cauchy slice. In the third line above we have used the localisation properties to commute $C$ with $\Pi[f_{+}]$ and $\Phi_{\textsc{a}}$ with $\Pi[f_{-}]$ (see figure4). The fourth line follows from the fact that the argument of $\Phi_{\textsc{a}}$, integrated over $x$, is proportional to the identity i.e. using Eq (58) which is shown in the appendix VII.5. Importantly, this leads to $\Phi_{\textsc{a}}$ dropping out from the expression, and as a result the details of A’s preparation disappear from C’s result.
We also see that in the last line of Eq (V.1) only the Kraus operators that are supported in $f_{-}$ (and not $f_{+}$) survive. Nevertheless, this does not have the form of an operation $\Phi_{\textsc{b}}[f_{-}]$, because of the double integration over $s$ and $s^{\prime}$ (note that the prefactor Eq (59) is generally different than $\delta(s-s^{\prime})$ for physical states of the pointer). Yet, we can rearrange things to show that
$\int$ $\mathrm{d}s\mathrm{d}s^{\prime}F_{\psi_{+}}(s-s^{\prime})\Pi^{\psi_{-}}_{s}[f_{-}]^{\dagger}C\Pi^{\psi_{-}}_{s^{\prime}}[f_{-}]$
$=\int\mathrm{d}q\Pi_{q}^{\psi}[f_{-}]^{\dagger}C\Pi_{q}^{\psi}[f_{-}],$ (61)
as we explicitly demonstrate in appendix VII.6.
This motivates the notion of effective Kraus operators (see next section). Overall, combining Eq (V.1) and Eq (V.1) we have shown that
$\text{tr}\left(\Phi_{\textsc{a}}\circ\Phi_{\textsc{b}}[f](C)\rho\right)=\text{tr}\left(\Phi_{\textsc{b}}[f_{-}](C)\rho\right).$ (62)
This is a no-signalling condition for B, since it is expected to hold for all maps $\Phi_{\textsc{a}}$ defined above, and all observables $C$ and states $\rho$. The properties of the local scattering map that corresponds to B ensure that only $\Phi_{\textsc{b}}[f_{-}]$ will affect the expectation value of any observable $C$ in the causal future.
V.2 Effective Kraus operators
In the previous subsection we proved (62), namely that the expectation values of $C$ do not depend on $\Phi_{\textsc{a}}$ and the details of B’s operation over the support of $f_{+}$. Yet, notice that we could have used any spacelike hypersurface for splitting $f$ as $f_{+}+f_{-}$. For any splitting, there are some points of $f_{-}$ that are spacelike to the future region R (see figure 7) which can be excluded by using another hypersurface that is ‘closer’ to the backward lightcone of R. This suggest that we can use covariance to get a stronger property of the Kraus operators, namely that they can be chosen to depend only on operators in the causal past of the region the algebra they act on is localised to, in a suitable limit. This holds for general regions, so observables localised to R will be updated by effective Kraus operators depending only on causally available data, independent of the hypersurface splitting we have used above.
Heuristically, the limit can be defined as follows: take a continuous family of hypersurfaces $\Sigma_{t}$ that approach the boundary of $J_{-}(\text{R})$ as $t\rightarrow\infty$, in such a way that they are all in the future of $J_{-}(\text{R})$ at all points and $\Sigma_{0}$ coincides in $\operatorname{supp}{f}$ with the hypersurface picked for the splitting into $f_{+},f_{-}$, see figure 7. Then let $\chi_{t}$ be the characteristic function of $J_{-}(\Sigma_{t})$, and $f_{t}(x)=f(x)\chi_{t}(x)$, and hence $f_{0}=f_{-}$, $\operatorname{supp}{f_{t}}\subset\operatorname{supp}{f_{t^{\prime}}}$ if $t>t^{\prime}$. It is clear that $f_{t}\rightarrow f\chi_{J_{-}(R)}$ as $t\rightarrow\infty$.
Finally, we use that the map $f\mapsto\Pi[f]$ is strongly continuous [oeckl2024spectraldecompositionfieldoperators]171717technically, this was proven only for the map between solutions and Kraus operators with $\psi$ Gaussian (see discussion of [oeckl2024spectraldecompositionfieldoperators] in section IV.2)., to see that observables localised to R are updated by the ‘effective’ would-be operators
$\Pi_{q}^{\psi}[f_{\textsc{r}}],\;\;f_{\textsc{r}}(x)=\begin{cases}f&x\in J_{-}(\text{R})\\ 0&x\notin J_{-}(\text{R}),\end{cases}$ (63)
depend only on the information causally available to R, and are independent of the exact choice of splitting. Note that these operators still correspond to the same measurement on the probe, in the sense that $\Pi_{q}^{\psi}[f],\Pi_{q}^{\psi}[f_{-}]$ are both conditional on the same value of the probe state $q$, and agree when acting on $\mathfrak{A}(\text{R})$. The reason that a limiting procedure is necessary, is that once cannot directly define (63) by ‘chopping’ the operation along the past lightcone (which introduces divergences), so (63) can only be understood in the limit. If this limit is well-defined, the effective update map is
$E_{f}(O)=\int\mathrm{d}q\Pi_{q}^{\psi}[f_{\textsc{r}}]^{\dagger}O\Pi_{q}^{\psi}[f_{\textsc{r}}]$ (64)
where we have that
$E_{f}(O)=$ $\int\mathrm{d}q\Pi_{q}^{\psi}[f_{\textsc{r}}]^{\dagger}O\Pi_{q}^{\psi}[f_{\textsc{r}}]$
$=O$ $\int\mathrm{d}q\Pi_{q}^{\psi}[f_{\textsc{r}}]^{\dagger}\Pi_{q}^{\psi}[f_{\textsc{r}}]=O$ (65)
whenever $O\in\mathfrak{A}(\text{R})$ with R being in the in-region of suppff, i.e. not intersecting its causal future.
Figure 7: A pictorial representation of the argument at the end of section V. The family of Cauchy surfaces $\Sigma_{t}$ can be seen to tend towards the intersection of the past null cone of R and $\operatorname{supp}{f}$.
Note that the suggested notion of effective operation, relies on covariance, i.e. freedom to split on any spacelike Cauchy slice, and on being able to take the limit that we sketched above. This is to guarantee that, in a sense, there are no superluminal processes within the support of $f$. Importantly, resolving Sorkin’s problem does not rely on this limiting procedure, but on the fact that there is always a choice of hypersurface such that $f_{-}$ will be fully spacelike to A, and so B cannot enable a superuminal signal from A to C (see figure 4). After all, even in the case when continuous additivity is not available, impossible operations are excluded by the Hammerstein property, as shown in the appendix VII.1.
V.3 Generality of the results
The property in Eq (57) is key in our anaylsis of causality and a motivation for effective Kraus operators. It it thus natural that we investigate when this condition holds in the local $S$-matrix formalism. In this section, we wish to understand what assumptions are required for identity (57) to hold more generally in the formalism of local $S$-matrices obeying
$S[f]=S[f_{+}+f_{-}]=S[f_{+}]S[f_{-}].$ (66)
We will assume that the $S$-matrix takes the form of a quantum controlled unitary. A quantum controlled unitary can be seen as first reading the value of the control system (in our case the probe) and then applying a unitary based on that control value to the other system (here the field). This is a direct generalisation of controlled gates in quantum circuits.
$|.S[f]=\int\mathrm{d}p\,U_{p}[f]\otimes\ket{p}\bra{p}$. (67)
The implication of continuous additivity (66) for (67) is that
$U_{p}[f_{+}]U_{p}[f_{-}]=U_{p}[f].$ (68)
This tells us that local $S$-matrices that implement quantum control factorise into matrices that factorise pointwise, which implies that
$\int\mathrm{d}s$ $\Pi^{\psi_{+}}_{q-s}[f_{+}]\Pi^{\psi_{-}}_{s}[f_{-}]$
$=$ $\int\mathrm{d}s\braket{q-s|S[f_{+}]|\psi_{+}}\braket{s|S[f_{-}]|\psi_{-}}$
$=\frac{1}{2\pi}$ $\int\mathrm{d}s\mathrm{d}p\mathrm{d}qU_{p}[f_{+}]U_{q}[f_{-}]\tilde{\psi}_{+}(p)\tilde{\psi}_{-}(q)e^{ipq}e^{is(p-q)}$
$=$ $\int\mathrm{d}pU_{p}[f_{+}]U_{p}[f_{-}]\tilde{\psi}_{+}(p)\tilde{\psi}_{-}(p)e^{ipq}$
$=$ $\int\mathrm{d}pU_{p}[f]\tilde{\psi}(p)=\Pi^{\psi}_{q}[f]$ (69)
where the penultimate equality comes from the usual convolution theorem, and the final equality comes from the assumption of continuous additivity, Eq (68), of $U_{p}[f]$. It should be noted that quantum control is not an assumption that holds in general.
For Kraus families consisting of a single element, the convolution identity takes the form of a Weyl identity:
$e^{iH_{+}}e^{iH_{-}}=e^{iH}e^{i\theta},$ (70)
where $H^{\pm},H$ are localised in the obvious sense. This is clearly satisfied when $H=a+\phi(f),a\mathbb{R}$, and fails for more general polynomials of fields: given a generic Cauchy slice $\Sigma$
$e^{i\phi(f)^{2}}\neq e^{iH_{+}}e^{iH_{-}}e^{i\theta}$ (71)
for $H^{\pm}$ localised in $\operatorname{supp}{f}\cap J^{\pm}(\Sigma)$. This is expected as generically only polynomials of first order in the fields lead to no-signalling unitary kicks. More details on this will be provided in [simmons_brukner_in_prep].
To summarise, we have shown that for some forms of $S$-matrix, i.e. quantum controlled, the Kraus operators directly inherit the factorisation property in the form of a convolution type identity as in Eq (57). Quantum controlled unitaries of this form are a very restricted set of evolutions, specifically ones where the Hamiltonian of the probe commutes with the interaction. For two fields coupled together, this is not a realistic assumption. While Eq (57) cannot be expected to hold in general, no-signalling will whenever $S$-matrices factorise. Further, eq (62) will also hold generally, whenever the $S$-matrix factorises, as
$S[f]^{\dagger}\mathfrak{A}(\text{C})S[f]=$ $S[f_{-}]^{\dagger}S[f_{+}]^{\dagger}\mathfrak{A}(\text{C})S[f_{+}]S[f_{-}]$
$=$ $S[f_{-}]^{\dagger}\mathfrak{A}(\text{C})S[f_{-}]$ (72)
implies $\Phi[f](C)=\Phi[f_{-}](C)$. Hence, while the convolution identity we used appears to be overly restrictive, the conclusions we made from it 1) no signalling and 2) the appearance of effective Kraus operators, hold far more generally. This implies that the conclusions we have made using the convolution identity are solid, and mostly independent of the $S$-matrix’s form, relying only on it factorising.
V.4 Measuring observables beyond $\phi(f)$
We have shown that there exists a very simple solvable model that can measure $\phi(f)$ for any $f$, in a completely causal way, and which recovers the results of [PhysRevD.105.025003, albertini2023ideal, oeckl2024spectraldecompositionfieldoperators], specifically that the Kraus operators should be given by a suitable $L^{2}$-normalised function of $\phi(f)$ (in the spectral calculus sense), which in our case depends on the choice of the state of the probe. However, $\phi(f)$ is a special operator in the algebra: it is unique in that products of smeared fields will generally not generate non-signalling unitary kicks. Since the non-signalling of a Gaussian measurement of $\phi(f)$ can be shown to follow from the non-signalling of unitary kicks by $e^{i\phi(f)}$ [PhysRevD.105.025003], it would be surprising if Gaussian (and more general $L^{2}$-normalised) measurements of powers of smeared fields were non-signalling. Indeed, it has been shown for a large class of products of smeared fields, that Gaussian measurements are signalling [PhysRevD.105.025003, albertini2023ideal].
One way to see this is to note that for sufficiently temporally thin $f$ and in a lattice approximation, $\phi(f)$ can be approximated as
$\phi(f)\sim A_{1}\otimes 1\dots\otimes 1+1\otimes A_{2}\dots\otimes 1$
$+\dots 1\otimes 1\dots\otimes A_{N}.$ (73)
Compact operators of this form can be measured in a non-signalling way [Borsten:2019cpc] with an update map that projects onto its eigenstate i.e. a non-destructive measurement. However, it is well known from quantum information [PhysRevLett.90.010402, PhysRevA.49.4331] that compact ‘non-local’ operators cannot be measured in a way that does not erase all local information. Non-local here refers to the fact that they cannot be written in a way that is completely local with respect to the tensor product decomposition e.g. $A\otimes B+A^{\prime}\otimes B^{\prime}\neq\tilde{A}\otimes 1+1\otimes\tilde{B}$. In field theory,
$F[\phi]=\int\mathrm{d}\mathsf{x}_{1}\dots\mathrm{d}\mathsf{x}_{n}F(\mathsf{x}_{1},\dots\mathsf{x}_{n})\phi(\mathsf{x}_{1})\dots\phi(\mathsf{x}_{n})$ (74)
is local if
$F(\mathsf{x}_{1},\dots\mathsf{x}_{n})=f(\mathsf{x}_{1})\delta(\mathsf{x}_{1}-\mathsf{x}_{2})\dots\delta(\mathsf{x}_{1}-\mathsf{x}_{n})$ (75)
[…]
This leads us to conclude that there are two distinct consequences of causality for measurements in QFT: 1) dynamical assumptions preclude the approximations used in the usual von Neumann derivation of ideal measurements, 2) non-local operators must be measured in such a way that the selective state update is destructive [PhysRevA.49.4331, PhysRevLett.90.010402]. While a projective measurement or even the measurement model we have studied here has Kraus operators that are simple spectral functions of the observable, and so final state is closely related to the initial state, for non-local observables, we expect the selective version of any causal update rule will lead to a final state that is very different from the original, destroying the initial state.
VI Conclusions
The impossibility of importing many ‘standard’ channels into QFT leaves a large gap in its foundations. Inspired by this, we have studied the relationship between non-signalling channels and unitaries using the local $S$-matrix approach, and related it to non-signalling measurement channels in QFT. Within this approach there are natural conditions, namely a hierarchy of factorisation conditions, which are sufficient for ruling out Sorkin’s impossible operations. Further, we have studied the operational implications for the resulting channels and Kraus operators, leading to causality conditions and the notion of effective Kraus operators. This provides a bridge between the abstract channels proposed in [PhysRevD.105.025003, albertini2023ideal, oeckl2024spectraldecompositionfieldoperators] and the rigorous FV framework [fewster2020quantum, PhysRevD.103.025017, fewster_asymptotic_2023].
The results in this paper hinged on combining quantum information notions of Stinespring and Kraus representations, and abstract notions of local QFT channels, which we introduced section II. With these notions, we reviewed examples of seemingly reasonable local channels that nevertheless lead to superluminal signalling, as typified by Sorkin’s impossible operations. In section II.3.1, the fact that $e^{i\phi(f)}$ does not signal while $e^{i\phi(f)^{2}}$ does, leads us to the notion of factorisation for local $S$-matrices in section III. We showed that local $S$-matrices obeying continuous additivity or the Hammerstein factorisation property cannot signal. With this general framework, we studied an explicit example of a local causally factorisable $S$-matrix $S[f]$, given by a von Neumann measurement of $\phi(f)$ in section IV.2, which is exactly solvable, and lead to an unsharp measurement of $\phi(f)$. Finally, we studied the properties of the $S$-matrix and the induced Kraus operators on the field in section V, and showed that, if continuous additivity holds, the Kraus operators can be chosen to be localised to a subregion arbitrarily close to being only in the causal past of the region they are updating, which we call effective Kraus operators. In principle this property appears stronger than what is required to rule out Sorkin’s impossible measurements, however it naturally follows from the local $S$-matrix formalism, and so it is expected to be a property of many physically implementable channels.
While the $S$-matrix framework we highlight is very general, the application to even the simplest model, of a trivially evolving pointer, has been fruitful. As well as providing a purification of abstract channels, it naturally includes dynamical information that does not appear in those channels constructed by hand. Specifically, we see a lower bound on the sharpness of a measurement of $\phi(f)$ given by $\Delta_{R}(f,f)^{\frac{1}{2}}$, and thus explicit dependence on the causal connectivity of the coupling region. To reach a sharp measurement, one must replace $f\mapsto f\lambda$, and take $\lambda\rightarrow 0$, amounting to changing the physical coupling (see e.g. [mandrysch2024quantumfieldmeasurementsfewsterverch]), and while also taking the probe state to be more sharp in $Q$. Similarly, we find that our model exhibits a generalisation of the convolution identity for Kraus operators found in [oeckl2024spectraldecompositionfieldoperators]. We’ve argued that this identity is a special case of a more general property of Kraus operators, and follows from the quantum controlled nature of the $S$-matrix in the simple model.
The connection between factorisation (continuous additivity and Hammerstein) and ruling out impossible measurements is not surprising, as they both address notions of causality. However, it does provide a direct link between an operational notion of no signalling in Sorkin scenarios, and a QFT notion of causality, i.e. factorisation of the $S$-matrix across any Cauchy surface. This stands in contrast with the conditions proposed in [sorkin1993impossible, Borsten:2019cpc, PhysRevD.105.025003, albertini2023ideal, Fuksa], which are much more global, in the sense that they depend non-trivially on the whole measurement scenario, and do not necessarily reduce to a condition that must hold locally within a given region, independently of the global scenario. Since the stronger form of continuous additivity implies microcausality of the Lagrangian, and microcausality of spacelike Heisenberg fields, we see a restoration of the connection between microcausality and no-signalling, which fits nicely into the AQFT framework.
It is tempting at this point to comment on the possible connections to a causal Stinespring’s theorem. Of course, the Stinespring dilation theorem is a general fact about unital CP channels between $C^{*}$-algebras, and so holds in QFT abstractly. However, in the literature a causal Stinespring theorem has been eluded to several times [PhysRevD.105.025003, fewster_asymptotic_2023], well summarised by Jubb “It would be useful to construct an explicit dictionary between update maps and specific probe models, and determine if this is possible in general. The latter would be analogous to Stinespring’s Dilation theorem, but with the added restriction of locality and causality on the unitary”. Relatedly, in this paper we have shown that the abstract maps of [PhysRevD.105.025003, albertini2023ideal, oeckl2024spectraldecompositionfieldoperators], when interpreted as functionals of the function $f$, can be dilated to an $S$-matrix $S[f]$ that obeys factorisation. In this case we found that the Kraus operators have general compositional properties that are directly related to the factorisation of $S[f]$. This has the flavour of a causal Stinespring theorem, or at least can be viewed as a hint for relativistic constraints on dilated channels. Since we have only focused on a small subset of local $S$-matrices when discussing the properties of the induced Kraus operators, we cannot make general statements. However, we believe that this analysis provides hints towards a causal Stinespring theorem, something we will present general results on elsewhere [simmons_brukner_in_prep].
Acknowledgements.
This research was funded in whole or in part by the Austrian Science Fund (FWF) [10.55776/F71] and [10.55776/COE1]. For open access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. This publication was made possible through the financial support of WOST (WithOutSpaceTime) grant from the John Templeton Foundation. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation. The authors are grateful to Robert Oeckl and José de Ramon Rivera for detailed feedback during this work. We also thank Jan Mandrysch and Miguel Navascués for discussions on the connections between our models. We are also grateful to Francisco Calderón and Doreen Fraser for insightful comments on earlier versions of this draft.
VII Appendix
VII.1 All operations with the Hammerstein property are possible
In this section we show that operations induced by local $S$-matrices obeying the first three conditions of section III.1 and the Hammerstein property [rejzner2019localitycausalityperturbativeaqft],
$S[f_{1}+f_{2}+f_{3}]=S[f_{1}+f_{2}]S[f_{2}]^{-1}S[f_{2}+f_{3}],$ (76)
[…]
$\text{tr}(\rho\mathcal{E}_{\textsc{a}}(S[f]^{\dagger}CS[f]))-\text{tr}(\rho S[f]^{\dagger}CS[f])=\delta,$ (77)
where $\mathcal{E}_{\textsc{a}}$ is local to A and $C\in\mathfrak{A}(\text{C})$. If $\delta=0$ for all $\rho,C$ and all Sorkin scenarios, then $S[f]$ is non-signalling.
[…]
By definition, there exists a Cauchy surface $\Sigma$ such that $J^{+}(\Sigma)\cap\text{B}=\text{B}_{+}$ is spacelike to C and $J^{-}(\Sigma)\cap\text{B}=\text{B}_{-}$ is spacelike to A. Also consider the following two regions $\tilde{\text{B}}_{+}=J^{+}(\text{A})\cap\text{B}$ and $\tilde{\text{B}}_{-}=J^{-}(\text{C})\cap\text{B})$. Note that $\tilde{\text{B}}_{\pm}\subsetneq\text{B}_{\pm}$, see figure8. Then we define the following (non-unique) continuous functions $f_{\pm}$
$f_{\pm}|_{\text{B}_{\mp}}=0,\;\;f_{\pm}|_{\tilde{\text{B}}_{\pm}}=f,\;\;f_{\pm}|_{\Sigma}=0$ (78)
i.e. $f_{\pm}$ vanishes in the future (past) of $\Sigma$, and coincides with $f$ in the future of A (past of C). Finally, let
$f_{0}=f-f_{+}-f_{-}.$ (79)
[…]
$\text{tr}(\rho\mathcal{E}_{\textsc{a}}(S[f]^{\dagger}CS[f]))=$ $\text{tr}(\rho\mathcal{E}_{\textsc{a}}(S[f_{-}+f_{0}]^{\dagger}S[f_{0}]S[f_{+}+f_{0}]^{\dagger}CS[f_{+}+f_{0}]S[f_{0}]^{\dagger}S[f_{-}+f_{0}]))$
$=$ $\text{tr}(\rho\mathcal{E}_{\textsc{a}}(S[f_{-}+f_{0}]^{\dagger}S[f_{0}]CS[f_{0}]^{\dagger}S[f_{-}+f_{0}]))$
$=$ $\text{tr}(\rho\mathcal{E}_{\textsc{a}}(S[f_{-}+f_{0}]^{\dagger}CS[f_{-}+f_{0}]))$
$=$ $\text{tr}(\rho(S[f_{-}+f_{0}]^{\dagger}CS[f_{-}+f_{0}]))$
$=$ $\text{tr}(\rho(S[f_{-}+f_{0}]^{\dagger}S[f_{0}]S[f_{+}+f_{0}]^{\dagger}CS[f_{+}+f_{0}]S[f_{0}]^{\dagger}S[f_{-}+f_{0}]))$
$=$ $\text{tr}(\rho(S[f]^{\dagger}CS[f])).$ (80)
To reach the second and third lines, we have used that $S[f_{+}+f_{0}]$ and $S[f_{0}]$ commute with $\mathfrak{A}(\text{C})$ respectively. To reach the fourth line, we have used that $S[f_{-}+f_{0}]^{\dagger}CS[f_{-}+f_{0}]$ is a product of operators that are localised spacelike to A, so $\mathcal{E}_{\textsc{a}}$ acts trivially by locality. Finally, we have used unitarity of the $S$-matrices and cyclicity of the trace to reach the final line. This implies $\delta=0$, and since we have not made any assumptions about the regions, or $\rho,C$, we can conclude that $S[f]$ is possible if it satisfies the Hammerstein property. Specifically, we can see that $\delta=0$ for all conditions implies $S[f]^{\dagger}\mathfrak{A}(\text{C})S[f],\mathfrak{A}(\text{A})]=0$ if $\mathfrak{A}(\text{C}),\mathfrak{A}(\text{A})]=0$, which implies that it is a non-signalling evolution [Borsten:2019cpc].
VII.2 Solving linear pointer coupling
Given the quadratic coupling specified by the term
$\int\mathrm{d}\text{vol }\mathcal{L}_{f}=\phi(f)\otimes P,\;\;\mathcal{L}_{f}=f(\mathsf{x})\phi(\mathsf{x})\otimes P,$ (81)
our $S$-matrix takes the form
$|.S[f]=\mathcal{T}e^{-i\phi(f)\otimes P}=\int dp\mathcal{T}e^{-ip\phi(f)}\otimes\ket{p}\bra{p}$. (82)
The Magnus expansion allows us to rewrite time ordered exponentials $\mathcal{T}\text{exp}(\cdot)=\text{exp}\sum_{n}\Xi_{n}(\cdot)$. Since the the pointer evolves trivially, the Lagrangian term $\mathcal{L}(x)=\phi(x)\otimes P$ implies
$\mathcal{L}(x),\mathcal{L}(x^{\prime})]=c1,$ (83)
and so only the terms that include at most one commutator survive,
$\Xi_{1}=-ip\phi(f),\;\Xi_{2}=-\frac{ip^{2}}{2}\Delta_{\textsc{r}}(f,f),\;\Xi_{n}=0,\;n>2.$ (84)
Finally, we have
[…]
VII.3 Continuous additivity of the local scattering maps in the linear pointer coupling
We wish to show that (85) satisfies
$S[f]=S[f_{+}]S[f_{-}],$ (86)
where $\in C_{0}(\mathcal{M}), f+=f$$_{+}=f\Theta,f_{-}=(1-\Theta)f$ where $\Theta(x)=\Theta(t)$ for some foliation $\mathcal{M}\cong\mathbb{R}_{t}\times\mathbb{R}^{3}$. In order to see this, we use the definitions $\phi(f),\phi(g)]=i\Delta(f,g)$ where
$\Delta(f,g)=\Delta_{\textsc{a}}(f,g)-\Delta_{\textsc{r}}(f,g)$ (87)
where $\Delta_{\textsc{a},\textsc{r}}$ the advanced and the retarded Green’s functions respectively. Note that
$\Delta(f_{+}\!,f_{-})=-\Delta(f_{-}\!,f_{+})=-\Delta_{\textsc{r}}(f_{-}\!,f_{+})$ (88)
and
$\Delta_{\textsc{r}}(f^{+}\mkern-5.0mu+\mkern-3.0muf^{-}\mkern-5.0mu,f^{+}\mkern-5.0mu+\mkern-3.0muf^{-})=\Delta_{\textsc{r}}(f_{+}\!,f_{+})+\Delta_{\textsc{r}}(f_{-}\!,f_{-})+\Delta_{\textsc{r}}(f_{-}\!,f_{+})$ (89)
since $\Delta_{\textsc{a}}(f_{-}\!,f_{+})=0$. Then,
[…]
$=S[f_{+}\mkern-5.0mu+\mkern-3.0muf_{-}]$ (90)
where the 2nd line comes from BCH
$e^{i\phi(f)}=e^{i\phi(f_{+})}e^{i\phi(f_{-})}e^{-i\Delta(f_{+},f_{-})/2},$ (91)
and 3rd from the propagator identities (88) and (89) above.
VII.4 Non-signalling convolution property of the Kraus operators in the linear pointer coupling
Suppose $\psi$ is a normalised $L^{2}(\mathbb{R})$ function, then the Kraus operators for a field coupled to a probe in state $\braket{x|\psi}=\psi(x)$, using the decomposition (85), are given by
$\Pi^{\psi}_{q}[f]$ $=\langle x|S[f]|\psi\rangle$
[…]
where $\tilde{\psi}$ is the Fourier transform of $\psi$. This simple model satisfies the following convolution property, which was been proven to be relevant for ensuring no-signalling [oeckl2024spectraldecompositionfieldoperators]
$\Pi^{\psi}_{q}[f]=\int\mathrm{d}s\,\Pi^{\psi_{+}}_{q-s}[f_{+}]\Pi^{\psi_{-}}_{s}[f_{-}],$ (93)
where $f_{\pm}$ are defined in the previous subsection, and $\psi_{\pm}$ are an arbitrary decomposition of the initial state of the probe $\psi$. To see this, we start from
[…]
Performing the $s$ and then $q$ integrals and using the BCH (91)
$\int\mathrm{d}s\,$ $\Pi^{\psi_{+}}_{q-s}[f_{+}]\Pi^{\psi_{-}}_{s}[f_{-}]$
[…]
$=\Pi^{\psi_{+}\star\psi_{-}}_{q}[f_{+}\!+f_{-}]$ (95)
where
$\widetilde{\psi_{+}\star\psi_{-}}(p)=\frac{1}{2\pi}\widetilde{\psi}(p).$ (96)
VII.5 A useful identity for the Kraus operators
Another identity of the Kraus operators that we need for our derivation of non-signalling is the following
[…]
To see this,
$\int$ $\mathrm{d}q\Pi^{\psi}_{q-s}[f]^{\dagger}\Pi^{\psi}_{q-s^{\prime}}[f]$
[…]
Note that for $s=s^{\prime}$ and for $\langle\psi|\psi\rangle=1$ this reduces to the usual trace preserving condition for Kraus operators,
[…]
VII.6 ‘Half’ an operation is still an operation
Here we show the equality in Eq (V.1). Consider $C$ local to a region that is spacelike to the support of $f_{+}$. Undoing the last step of (V.1), what we want to show is that
$\int\mathrm{d}q\mathrm{d}s\mathrm{d}s^{\prime}\text{tr}\left(C\Pi^{\psi_{-}}_{s^{\prime}}[f_{-}]\rho\Pi^{\psi_{-}}_{s}[f_{-}]^{\dagger}\left(\Pi^{\psi_{+}}_{q-s}[f_{+}]^{\dagger}\Pi^{\psi_{+}}_{q-s^{\prime}}[f_{+}]\right)\right)=\int\mathrm{d}q\text{tr}\left(\Pi_{q}^{\psi}[f_{-}]^{\dagger}C\Pi_{q}^{\psi}[f_{-}]\rho\right)$ (99)
In order to do so, we expand the left hand side as follows
$\int\mathrm{d}q\mathrm{d}s\mathrm{d}s^{\prime}\text{tr}\left(C\Pi^{\psi_{-}}_{s^{\prime}}[f_{-}]\rho\Pi^{\psi_{-}}_{s}[f_{-}]^{\dagger}\left(\Pi^{\psi_{+}}_{q-s}[f_{+}]^{\dagger}\Pi^{\psi_{+}}_{q-s^{\prime}}[f_{+}]\right)\right)=\frac{1}{(2\pi)^{2}}\int\mathrm{d}q\mathrm{d}s\mathrm{d}s^{\prime}\mathrm{d}p\mathrm{d}n\mathrm{d}k\mathrm{d}m\widetilde{\psi_{+}}^{*}(p)\widetilde{\psi_{-}}^{*}(n)\widetilde{\psi_{+}}(k)\widetilde{\psi_{-}}(m)$
$\times\text{exp}\left(\frac{i}{2}(\Delta_{\textsc{r}}(f_{+},f_{+})(p^{2}-k^{2})+\Delta_{\textsc{r}}(f_{-},f_{-})(n^{2}-m^{2})-ip(q-s)-iqs+ik(q-s^{\prime})+ims^{\prime}\right)$
$\times\text{tr}\left(Ce^{-im\phi[f_{-}]}\rho e^{in\phi[f_{-}]}e^{-i\phi[f_{+}](p-k)}\right)$ (100)
Performing the $s,s^{\prime}$ integrals, we get,
$\int\mathrm{d}q\mathrm{d}s\mathrm{d}s^{\prime}\text{tr}\left(C\Pi^{\psi_{-}}_{s^{\prime}}[f_{-}]\rho\Pi^{\psi_{-}}_{s}[f_{-}]^{\dagger}\left(\Pi^{\psi_{+}}_{q-s}[f_{+}]^{\dagger}\Pi^{\psi_{+}}_{q-s^{\prime}}[f_{+}]\right)\right)=\int\mathrm{d}q\mathrm{d}p\mathrm{d}k\widetilde{\psi_{+}}^{*}(p)\widetilde{\psi_{-}}^{*}(p)\widetilde{\psi_{+}}(k)\widetilde{\psi_{-}}(k)$
$\times\text{exp}\left(\frac{i}{2}(\Delta_{\textsc{r}}(f_{+},f_{+})+\Delta_{\textsc{r}}(f_{-},f_{-}))(p^{2}-k^{2})-iq(p-k)\right)\text{tr}\left(Ce^{-ik\phi[f_{-}]}\rho e^{ip\phi[f_{-}]}e^{-i\phi[f_{+}](p-k)}\right)$
$=\int\mathrm{d}p|\widetilde{\psi}(p)|\text{tr}\left(e^{ip\phi[f_{-}]}Ce^{-ip\phi[f_{-}]}\rho\right),$ (101)
where in the last step we performed the $x$ integral and then $k$ integral and used that $\widetilde{\psi_{+}\star\psi_{-}}=\frac{1}{2\pi}\widetilde{\psi}$. Note that the Dirac delta $\delta(k-p)$ ensures that the factors including $\phi(f_{+})$ disappear. Finally, to reintroduce the expression of the local Kraus operators (using the fact that the update map is not sensitive to any phase multiplying the Kraus operators) we perform the following trick
$\int\mathrm{d}p|\widetilde{\psi}(p)|\text{tr}\left(e^{ip\phi[f_{-}]}Ce^{-ip\phi[f_{-}]}\rho\right)=\frac{1}{2\pi}\int$ $\mathrm{d}q\mathrm{d}p\mathrm{d}k\widetilde{\psi}(p)^{*}\widetilde{\psi}(q)\text{exp}\left(\frac{i}{2}(p^{2}-k^{2})\Delta_{\textsc{r}}(f_{-},f_{-})-iq(p-k)\right)$
$\times\text{tr}\left(e^{ip\phi[f_{-}]}Ce^{-ik\phi[f_{-}]}\rho\right)=\int\mathrm{d}q\text{tr}\left(\Pi_{q}^{\psi}[f_{-}]^{\dagger}C\Pi_{q}^{\psi}[f_{-}]\rho\right)$ (102)
as claimed.

### v2

@@
 I Introduction
-The topic of impossible measurements [sorkin1993impossible] has received renewed attention over the past few years.
-A multitude of approaches have been developed to address the impossible measurement scenarios (see [papageorgiou2024eliminating] and references therein).
+The topic of impossible measurements [47] has received renewed attention over the past few years.
+A multitude of approaches have been developed to address the impossible measurement scenarios (see [38] and references therein).
 The clear moral that has emerged is that the set of quantum mechanical operations, or channels, typically allowed in quantum information is far too large when brought to bear on relativistic theories.
@@
 In fact, the established empirical success of QFT relies on the asymptotic scattering paradigm, which is not suitable for describing local operations, i.e.
-operations over bounded regions of spacetime (for historical context see [Fraser2023]).
-Sorkin’s examples of impossible operations [sorkin1993impossible], and recent refinements [Borsten:2019cpc], showed that operations built out of local operators (i.e.
+operations over bounded regions of spacetime (for historical context see [19]).
+Sorkin’s examples of impossible operations [47], and recent refinements [6], showed that operations built out of local operators (i.e.
 local Kraus operators) generally allow superluminal signalling.
-Specifically, this includes ideal measurements of many important observables [Benincasa_2014, albertini2023ideal, PhysRevD.105.025003], leading to a number of papers suggesting non-ideal measurement channels, both abstract [PhysRevD.105.025003, Gisin2024towardsmeasurement, albertini2023ideal, oeckl2024spectraldecompositionfieldoperators] and concrete i.e.
-derived from some purified model [mandrysch2024quantumfieldmeasurementsfewsterverch].
+Specifically, this includes ideal measurements of many important observables [2, 1, 28], leading to a number of papers suggesting non-ideal measurement channels, both abstract [28, 22, 1, 35] and concrete i.e.
+derived from some purified model [30].
 The former use a number of no-signalling conditions on the abstract channels or their Kraus operators, while the latter rely on properties of the purified dynamics.
-From the opposite direction, a rigorous framework for quantum measurement theory in algebraic QFT has been proposed by Fewster and Verch (FV framework)[fewster2020quantum, PhysRevD.103.025017, Fewster2023], based on purified dynamics.
+From the opposite direction, a rigorous framework for quantum measurement theory in algebraic QFT has been proposed by Fewster and Verch (FV framework)[16, 7, 17], based on purified dynamics.
 Together, these results point towards a general theory of measurement and operations in QFT.
@@
 The main results of the paper are as follows.
-We adapt the local $S$-matrix formalism due to Bogoliubov [pittphilsci24053] to the setting of Sorkin’s impossible operations.
+We adapt the local $S$-matrix formalism due to Bogoliubov [4, 3] to the setting of Sorkin’s impossible operations.
 We demonstrate which of the factorisation conditions are necessary or sufficient for blocking impossible operations, faster than light signalling and retrocausality.
@@
 $\Pi_{q}^{\psi}[f]=\int\mathrm{d}s\,\Pi_{q-s}^{\psi_{+}}[f_{+}]\Pi_{s}^{\psi_{-}}[f_{-}].$ (2)
-Finally, we use an exactly solvable concrete measurement model that can seen as a relativistic von Neumann measurement model, where a local pointer variable is linearly coupled to a smeared scalar field operator, as $\phi(f)\otimes P$ [PhysRevD.109.065024].
+Finally, we use an exactly solvable concrete measurement model that can seen as a relativistic von Neumann measurement model, where a local pointer variable is linearly coupled to a smeared scalar field operator, as $\phi(f)\otimes P$ [37].
 We find that the $S$-matrices and Kraus operators obey the conditions (1) and (2), and so this forms a concrete example of the link between the two representations of channels and no-signalling conditions.
 Further, by solving the local $S$-matrix model we obtain an explicit class of measurement channels on the field whose sharpness is fundamentally limited by the retarded field propagation within the support of $f$, thereby imposing an intrinsic bound on the accuracy of local field measurements.
-These channels coincide with the abstract continuous measurement channels studied in [PhysRevD.105.025003, albertini2023ideal, oeckl2024spectraldecompositionfieldoperators], once this new sharpness constraint is taken into account.
+These channels coincide with the abstract continuous measurement channels studied in [28, 1, 35], once this new sharpness constraint is taken into account.
 The paper is structured as follows: In section II we briefly introduce standard measurement theory and a minimal (algebraic) QFT set-up.
-We review an abstract notion of operation in the QFT set-up, and briefly present the challenges posed by the impossible measurements scenario, using a simple example of impossible unitary operations introduced in [PhysRevD.105.025003].
+We review an abstract notion of operation in the QFT set-up, and briefly present the challenges posed by the impossible measurements scenario, using a simple example of impossible unitary operations introduced in [28].
 In section III we introduce the local $S$-matrix formalism, to discuss the properties of relativistically valid unitary operations in QFT that can (in principle) represent local (not asymptotic) interactions.
-We take the local $S$-matrix to represent the coupling of the field to a probe system [hellwigkraus1970b] and we examine the operational consequences of a hierarchy of factorisation properties of the $S$-matrix.
+We take the local $S$-matrix to represent the coupling of the field to a probe system [26] and we examine the operational consequences of a hierarchy of factorisation properties of the $S$-matrix.
 In section IV we explicitly introduce field operations induced by probes, and the induced Kraus operators in the field’s Hilbert space that are fixed by the local $S$-matrix and the characteristics of the probe system.
 We consider a simple model for measurements of $\phi(f)$, and consider the limitations to the accuracy of field measurements that come from the causal (retarded) field propagation within the support of $f$.
-We show how this model relates to the families of Kraus operators for field measurements introduced in Albertini and Jubb [albertini2023ideal] and Oeckl [oeckl2024spectraldecompositionfieldoperators].
+We show how this model relates to the families of Kraus operators for field measurements introduced in Albertini and Jubb [1] and Oeckl [35].
 Specifically, the result that the sharpness of a measurement of $\phi(f)$ is bounded from below by $\Delta_{\textsc{r}}(f,f)^{1/2}$, i.e.
@@
 These operations, when extended to QFT, can lead to signalling across spacelike separated regions.
-The problem of impossible operations, which was first introduced by Sorkin [sorkin1993impossible], was meant to highlight the friction between no-signalling and the quantum operations that are typically assumed in quantum mechanics.
-For examples see [Benincasa_2014] and more recently [PhysRevD.105.025003, albertini2023ideal].
+The problem of impossible operations, which was first introduced by Sorkin [47], was meant to highlight the friction between no-signalling and the quantum operations that are typically assumed in quantum mechanics.
+For examples see [2] and more recently [28, 1].
 In this section we briefly review the basics of standard quantum measurement theory, we introduce a minimal (algebraic) framework for QFT and we discuss the problem of ‘impossible measurements’.
@@
 II.1 Standard measurement theory
-Quantum mechanics has a standard measurement theory, see [busch_book_2016, busch_standard_1996].
+Quantum mechanics has a standard measurement theory, see [10, 11].
 Despite ongoing ambiguity regarding what precisely qualifies certain interactions as measurements – a central aspect of the quantum measurement problem – there exist standard prescriptions for extracting probabilistic predictions for any measurement or general operation defined within the theory.
@@
 II.2 Set-up
-Below we introduce the set-up and we review an example from [PhysRevD.105.025003] that is most relevant for the measurement model that we will consider in section IV.2.
+Below we introduce the set-up and we review an example from [28] that is most relevant for the measurement model that we will consider in section IV.2.
 We focus on the real scalar field.
@@
 It follows that any bounded function of $\phi(f)$, defined by the spectral calculus, is in $\mathfrak{A}(\text{R})$.
-For more details on this construction, see [fewster2019algebraic].
-Using the properties of the commutator, it follows, from strong commutativity of the fields [Borchers:1992rq], that spacelike algebras commute,
+For more details on this construction, see [18].
+Using the properties of the commutator, it follows, from strong commutativity of the fields [5], that spacelike algebras commute,
 $\mathfrak{A}(\text{R}),\mathfrak{A}(\text{R}^{\prime})]=0,$ (19)
@@
 More generally, in axiomatic approaches to QFT (e.g.
-algebraic QFT [haag_kastler_1964]) one postulates the axiom of microcausality, namely that spacelike separated algebras of observables commute.
-This is a kinematical ‘separability’ assumption that, roughly, splits the field degrees of freedom between causally disconnected parts [earman2014relativistic].
-Note that microcausality alone does not have an a priori operational consequence [Ruetsche2021, Calderon2022, soulas2025proof] e.g.
+algebraic QFT [23]) one postulates the axiom of microcausality, namely that spacelike separated algebras of observables commute.
+This is a kinematical ‘separability’ assumption that, roughly, splits the field degrees of freedom between causally disconnected parts [15].
+Note that microcausality alone does not have an a priori operational consequence [44, 12, 48] e.g.
 in terms of signalling.
-Once a notion of operation is introduced, microcausality alone is generally not sufficient for blocking superluminal signalling (see also [redei_how_2010]).
-Sorkin showed that [sorkin1993impossible], despite the microcausality assumption, even the most basic local operations in QFT, such as unitary kicks and ideal measurements – when followed by the Lüders post-measurement state update for non-selective measurements – can lead to superluminal signaling between spacelike separated regions (see also [Borsten:2019cpc, PhysRevD.105.025003, oeckl2024spectraldecompositionfieldoperators]).
+Once a notion of operation is introduced, microcausality alone is generally not sufficient for blocking superluminal signalling (see also [41]).
+Sorkin showed that [47], despite the microcausality assumption, even the most basic local operations in QFT, such as unitary kicks and ideal measurements – when followed by the Lüders post-measurement state update for non-selective measurements – can lead to superluminal signaling between spacelike separated regions (see also [6, 28, 35]).
 […]
@@
 For region A to not signal to region C, the expectation values of observables in C must not depend on the choice $\lambda$ of which operation $\Phi^{\lambda}_{\textsc{a}}$ is implemented in A.
-Namely, the condition for no impossible operations (which has previously been considered in [Borsten:2019cpc, deramon2021relativistic, PhysRevD.105.025003, oeckl2024spectraldecompositionfieldoperators]) is simply the following: let $\rho$ be the global state of the field and $C$ an observable over region C, then it should hold that
+Namely, the condition for no impossible operations (which has previously been considered in [6, 14, 28, 35]) is simply the following: let $\rho$ be the global state of the field and $C$ an observable over region C, then it should hold that
 $\text{tr}\left(\Phi_{\textsc{a}}^{\lambda}(\Phi_{\textsc{b}}(C))\rho\right)=\text{tr}\left(\Phi_{\textsc{b}}(C)\rho\right),\mbox{ }\forall\lambda.$ (21)
@@
 II.3.1 Example of signalling and non-signalling unitaries
-The most basic example of failure of this condition using unitary operators in QFT was given in [PhysRevD.105.025003] and uses the real scalar field construction given above. Letting
+The most basic example of failure of this condition using unitary operators in QFT was given in [28] and uses the real scalar field construction given above. Letting
 $\Phi^{\lambda}_{\textsc{a}}(\cdot)=e^{i\phi(h_{\lambda})}\cdot e^{-i\phi(h_{\lambda})},$
@@
 $\text{tr}\left(\Phi_{\textsc{a}}^{\lambda}(\Phi_{\textsc{b}}(C))\rho\right)=\Delta(f,g).$ (25)
-The failure of the no-signalling condition (21) in Eq (23) can be traced back to the fact that $\Phi_{\textsc{b}}(C)$ will not commute with $A$ in general, even though $[A,C]=0$ by microcausality of operators in AQFT [Borsten:2019cpc], while commutation does hold for Eq (25).
+The failure of the no-signalling condition (21) in Eq (23) can be traced back to the fact that $\Phi_{\textsc{b}}(C)$ will not commute with $A$ in general, even though $[A,C]=0$ by microcausality of operators in AQFT [6], while commutation does hold for Eq (25).
 Then, it is natural to suspect that the reason why $\Phi_{\textsc{b}}=e^{i\phi(f)^{2}}\cdot e^{-i\phi(f)^{2}}$ is problematic is that $\phi(f)^{2}$ is a non-local operator (given by a double integration over spacetime in this case).
@@
 Bogoliubov’s local $S$-matrix  [^4] The first attempt to formulate relativistic quantum theory in terms of the $S$-matrix is due to Heisenberg, who introduced a ‘global’ $S$-matrix satisfying unitarity and Lorentz invariance.
-The local $S$-matrix approach was later introduced by Stueckelberg and Bogoliubov, reintroducing local degrees of freedom and demanding additional causality conditions [pittphilsci24053].
-formalism [bogoliubov1960introduction] is a local and covariant approach to describing interactions in QFT, and so is a natural place for us to start.
-It plays a major role in perturbative algebraic QFT (see [rejzner2019localitycausalityperturbativeaqft] and references within).
+The local $S$-matrix approach was later introduced by Stueckelberg and Bogoliubov, reintroducing local degrees of freedom and demanding additional causality conditions [3].
+formalism [4] is a local and covariant approach to describing interactions in QFT, and so is a natural place for us to start.
+It plays a major role in perturbative algebraic QFT (see [42] and references within).
 A local $S$-matrix will be denoted as $S[f]$ where $f$ a compactly supported smearing function.
@@
 $\mathcal{L}_{I}(\mathsf{x})=-f_{\text{const}}\phi(\mathsf{x})^{4}$, and which map between the in and out Hilbert spaces.
-While the local $S$-matrix formalism was originally conceived, and still used in perturbative AQFT (pAQFT), to define scattering interactions in cases where the asymptotic $S$-matrix might not be well defined, it also provides a useful model for more general localised interactions, which may include scenarios beyond scattering.
+While the local $S$-matrix formalism was originally conceived, and still used in perturbative AQFT (pAQFT), to define scattering interactions in cases where the asymptotic $S$-matrix might not be well defined (for historical context see [20, 3]), it also provides a useful model for more general localised interactions, which may include scenarios beyond scattering.
 Further, as we will see, it comes endowed with a hierarchy of causal conditions which are very well suited for studying causality in finite time processes.
@@
 [^5] This condition can be imposed on a Dyson expansion of an $S$-matrix by assuming the unsmeared Lagrangian is microcausal.
-We call continuous additivity [bogoliubov1960introduction] the following property: given a hypersurface $\Sigma$ that intersects the region $\text{supp}f$ (see figure 3) we consider the decomposition $f=f_{+}+f_{-}$ where $\text{supp}f_{\pm}=\text{supp}f\cap J^{\pm}(\Sigma)$.
+We call continuous additivity [4] the following property: given a hypersurface $\Sigma$ that intersects the region $\text{supp}f$ (see figure 3) we consider the decomposition $f=f_{+}+f_{-}$ where $\text{supp}f_{\pm}=\text{supp}f\cap J^{\pm}(\Sigma)$.
 Then
 $S[f_{+}]S[f_{-}]=S[f_{+}+f_{-}].$ (29)
-This property allows us to decompose the dynamical evolution in past and future evolution with respect to any given foliation.
+This property allows us to decompose the dynamical evolution in past and future evolution with respect to any given foliation [^6] Note that by decomposing $f$ in this way, $f_{\pm}$ will generally not be smooth.
+Any mathematical problems that may arise from this (e.g.
+surface divergences, see [3, 20]) do not arise in the more complex Hammerstein factorisation, where all functions can be kept $C^{\infty}$ smooth.
+.
 Figure 3: Continuous additivity of an $S$-matrix corresponds to selection of a (partial) Cauchy surface $\Sigma$ that bisects the support of $f$.
@@
     $S[f]=S[f_{+}]S[f_{-}].$ (31)
-A few comments are in order regarding the meaning of Eq (31), since this property is central in our analysis of (retro)causality [^7] More abstractly, in local operational frameworks (see [oeckl_local_2019]) one demands a similar compositionality property of null probes (that represent dynamics) that share a boundary.
+A few comments are in order regarding the meaning of Eq (31), since this property is central in our analysis of (retro)causality [^8] More abstractly, in local operational frameworks (see [33]) one demands a similar compositionality property of null probes (that represent dynamics) that share a boundary.
 .
-For our purposes, we demand (31) inspired by the causal perturbation theory program888where the condition is normally introduced in order to help defining the local algebras of a perturbative and causal QFT, see [Ilyin:1978gq, Brunetti_2000, rejzner2019localitycausalityperturbativeaqft] and a recent project for non-perturbative AQFT [Buchholz_2020].
+For our purposes, we demand (31) inspired by the causal perturbation theory program999where the condition is normally introduced in order to help defining the local algebras of a perturbative and causal QFT, see [27, 8, 42] and a recent project for non-perturbative AQFT [9].
 In these works, it is analysed in detail how restrictive this assumption is for realistic interactions, which of course goes beyond our purposes., as well as weaker factorisation conditions (see section VII.1) to analyse their operationally meaningful consequences, in the context of signalling, in a simple measurement scenario involving a probe locally coupled to a field.
 A final comment about (31) is that it is key that we are cutting $f$ on a Cauchy slice, i.e.
-a spacelike hypersurface, rather than a lightlike hypersurface such as one that coincides with A or C’s lightcones, as factorising the local $S$-matrix on the lightcone would introduce divergences (which can be seen using microlocal analysis [Brunetti_2000]).
+a spacelike hypersurface, rather than a lightlike hypersurface such as one that coincides with A or C’s lightcones, as factorising the local $S$-matrix on the lightcone would introduce divergences (which can be seen using microlocal analysis [8]).
 Finally, so far our discussion has been limited to a QFT—whether consisting of multiple fields or a single one—in which the scattering matrix is restricted to act only within a single region (B).
@@
 which is a coupling between the field and one probe, localised to some region $R_{i}$.
-Approaches to measurement theory in QFT in which the measurement process is modelled as a local scattering process and operations are induced by probes, include the detector-based measurement theory [Polo_Gomez_2022] and the Fewster-Verch (FV) framework in algebraic QFT [fewster2020quantum].
+Approaches to measurement theory in QFT in which the measurement process is modelled as a local scattering process and operations are induced by probes, include the detector-based measurement theory [39] and the Fewster-Verch (FV) framework in algebraic QFT [16].
 In the detector-based approach, one works with (a general class) of Hamiltonian densities that model the detector-field coupling e.g.
@@
 Note that the approach we are taking here, is more abstract than the first (since it does not rely on particular interaction Hamiltonians and the Dyson expansion), but more concrete than the latter, since we assume the existence of these local scattering maps that implement the scatteting morphism in a Hilbert space.
-Partial or exact resolution to the impossible measurements scenario have been offered respectively in the detector-based approach [benincasa2014quantum, deramon2021relativistic] and in the FV-framework [PhysRevD.103.025017].
+Partial or exact resolution to the impossible measurements scenario have been offered respectively in the detector-based approach [2, 14] and in the FV-framework [7].
 In either case, an important ingredient of the resolution was an analogue causal factorisation property (32).
-An abstract version of causal factorisation is assumed to hold in the FV framework  [^10] The abstract version of the causal factorisation condition is that $\Theta_{\textsc{a+b}}=\Theta_{\textsc{b}}\circ\Theta_{\textsc{a}}$, which has to be proven for concrete models in this framework., while in [deramon2021relativistic] it was argued that that this property holds for a general class of Unruh-DeWitt-type detector models coupled locally to a coupled field over disjoint, causally orderable regions.
+An abstract version of causal factorisation is assumed to hold in the FV framework  [^11] The abstract version of the causal factorisation condition is that $\Theta_{\textsc{a+b}}=\Theta_{\textsc{b}}\circ\Theta_{\textsc{a}}$, which has to be proven for concrete models in this framework., while in [14] it was argued that that this property holds for a general class of Unruh-DeWitt-type detector models coupled locally to a coupled field over disjoint, causally orderable regions.
 Finally, all the measurement schemes considered in QFT correspond to so-called pre-measurement processes, involving only the unitary interaction between the field and the probe.
-Strictly speaking, the measurement is completed only once an outcome has been obtained, which would require, in turn, a measurement of the probe and so on, leading to an infinite regress, see [mandrysch2024quantumfieldmeasurementsfewsterverch].
+Strictly speaking, the measurement is completed only once an outcome has been obtained, which would require, in turn, a measurement of the probe and so on, leading to an infinite regress, see [30].
 In this work we ignore this final measurement stage and, following common jargon, refer to the pre-measurement stage itself as the “measurement process”.
@@
 More assumptions about the probe-field interactions in the coupling regions have to be made, i.e.
-the locality of the detector’s local current in the detector model approach [deramon2021relativistic]  [^11] In [deramon2021relativistic] the authors show that (seemingly local) Hamiltonian densities of the form $h(\mathsf{x})=\Lambda(\mathsf{x})J(\mathsf{x})\otimes\phi(\mathsf{x})$, where $J(\mathsf{x})$ is the current associated to a probe, can give rise to impossible measurements if the probe’s dynamical current is not microcausal (i.e.
-if $J(\mathsf{x}),J(\mathsf{y})]\neq 0$ for $\mathsf{x},\mathsf{y}$ is not spacelike separated) by using the Dyson expansion of $S=\mathcal{T}\text{exp}i\int h(\mathsf{x})\mathrm{d}\mathsf{x}$.
+the locality of the detector’s local current in the detector model approach [14]  [^12] In [14] the authors show that (seemingly local) Hamiltonian densities of the form $h(\mathsf{x})=\Lambda(\mathsf{x})J(\mathsf{x})\otimes\phi(\mathsf{x})$, where $J(\mathsf{x})$ is the current associated to a probe, can give rise to impossible measurements if the probe’s dynamical current is not microcausal (i.e.
+if $J(\mathsf{x}),J(\mathsf{y})]\neq 0$ for $\mathsf{x},\mathsf{y}$ spacelike separated) by using the Dyson expansion of $S=\mathcal{T}\text{exp}\,i\int h(\mathsf{x})\mathrm{d}\mathsf{x}$.
 , or the time-slice property of the coupled theory in the FV framework  [^13] The exact relation between continuous additivity and other dynamical axioms e.g.
 the time-slice property in the formal setting of (p)AQFT goes way beyond our purposes (for related results see e.g.
-[ruep_thesis_2022]).
-[PhysRevD.103.025017].
+[43]).
+[7].
 Our claim is that Eq (31) captures the properties of a causal probe-field coupling that arises from the additional assumptions needed in the concrete and abstract frameworks, while making it clear that this can be directly related to internal factorisation properties of the interactions.
@@
 We have introduced two notions of factorisations of $S$-matrices, causal factorisation and continuous additivity, that respectively correspond to composing disjoint interactions or decomposing a single interaction.
-Overall, there is a hierarchy of spacetime factorisation properties, as explained in [Brunetti_2000].
+Overall, there is a hierarchy of spacetime factorisation properties, as explained in [8].
 In this section, we will briefly outline them here for completeness, and generalise our proof of no impossible operations to a weaker condition.
@@
 For this reason, here we will also consider weaker factorisation conditions that are more widely assumed to hold for local interactions.
-One such weaker condition, called the Hammerstein property, is typically assumed in causal perturbation theory literature [rejzner2019localitycausalityperturbativeaqft], where it is taken as continuous additivity Eq (31) may be too strong for very singular interactions [Brunetti_2000].
+One such weaker condition, called the Hammerstein property, is typically assumed in causal perturbation theory literature [42], where it is taken as continuous additivity Eq (31) may be too strong for very singular interactions [8].
 The Hammerstein property is that
@@
 where we have used the stronger condition Eq (29) to factorise, and used that $S[0]=S[f]S[f]^{\dagger}\implies S[f]^{\dagger}=S[-f]$.
-Importantly, the Hammerstein property is also sufficient to rule out Sorkin problems, as we explicitly show in detail in the appendix VII.1 and sketch here.
+Importantly, the Hammerstein property is also sufficient to rule out Sorkin problems, as we explicitly show in detail in Appendix VII.1 and sketch here.
 Figure 5: The spacetime factorisation of $f$ into $f_{+}$ (red) $f_{0}$ (white) and $f_{-}$ (blue), where we note the disjoint and time ordered supports of $f_{\pm}$, and the support overlaps of $f_{0}$ with $f_{\pm}$.
@@
 This implies that $S[f_{+}+f_{0}]S[f_{0}]^{\dagger}$ commutes with $C$ and $S[f_{-}+f_{0}]^{\dagger}CS[f_{-}+f_{0}]$ is localised spacelike to A.
-Following a very similar derivation to Eq (III.1.1) demonstrates that there is no signalling in a Sorkin scenario where $\Phi_{\textsc{b}}$ is induced by an $S$-matrix satisfying the Hammerstein property  [^13] Resolutions of the Sorkin problem in the abstract algebraic setting of the FV framework rely on the time-slice property of the coupled theory [PhysRevD.103.025017, papageorgiou2024eliminating].
-At the level of the local $S$-matrix it has been shown that the time-slice property can be derived only from (36) and normalisation $S(0)=\mathsf{1}$ [chilian2009time], which explains why we were able to demonstrate the link between (36) and no impossible operations in appendix VII.1..
+Following a very similar derivation to Eq (III.1.1) demonstrates that there is no signalling in a Sorkin scenario where $\Phi_{\textsc{b}}$ is induced by an $S$-matrix satisfying the Hammerstein property  [^14] Resolutions of the Sorkin problem in the abstract algebraic setting of the FV framework rely on the time-slice property of the coupled theory [7, 38].
+At the level of the local $S$-matrix it has been shown that the time-slice property can be derived only from (36) and normalisation $S(0)=\openone [1$3], which explains why we were able to demonstrate the link between (36) and no impossible operations in Appendix VII.1..
 For the full details, see Appendix VII.1.
@@
 In the same vein, we will consider in this section the realisation of non-selective operations in QFT through interactions between the system (i.e., the field) and a probe.
-Then, a way to argue that unitary “kicks” which violate no-signalling are not physically realisable is to show that they cannot be implemented by any valid local interaction between the system and a probe (see also [muchverch]).
+Then, a way to argue that unitary “kicks” which violate no-signalling are not physically realisable is to show that they cannot be implemented by any valid local interaction between the system and a probe (see also [31]).
 Below, we define local Kraus operators as induced by a suitable local coupling $S[f]$ between the field and the probe, and we provide an example of a concrete model for field measurements.
@@
 The attempts of modelling local operations in QFT required a departure from the asymptotic scattering paradigm.
-In a series of papers [hellwigkraus1969, hellwigkraus1970b] Hellwig and Kraus model relativistic quantum measurement as a local scattering process.
-Their motivation for introducing the so-called Kraus operators, as induced through the dynamical coupling to some probe, was to explicate the operational interpretation of the local algebras that was implicit in the axiomatic approaches to QFT [haag_kastler_1964].
-Their measurement scheme can be viewed as predecessor of recently proposed measurement frameworks in QFT [Polo_Gomez_2022, fewster2020quantum] that involve some (abstract or concrete) representation of dynamics between system and probe (for historical context see [Fraser2023]).
+In a series of papers [25, 26] Hellwig and Kraus model relativistic quantum measurement as a local scattering process.
+Their motivation for introducing the so-called Kraus operators, as induced through the dynamical coupling to some probe, was to explicate the operational interpretation of the local algebras that was implicit in the axiomatic approaches to QFT [23].
+Their measurement scheme can be viewed as predecessor of recently proposed measurement frameworks in QFT [39, 16] that involve some (abstract or concrete) representation of dynamics between system and probe (for historical context see [19]).
 Consider the combined Hilbert space $\mathcal{H}^{(s)}\otimes\mathcal{H}^{(p)}$ of a system (field) with Hilbert space $\mathcal{H}^{(s)}$ and some probe with Hilbert space $\mathcal{H}^{(p)}$.
@@
 […]
-Kraus operators in scattering-based frameworks for local measurement in QFT have been derived previously in [SmithAlexander2017, Polo_Gomez_2022, PhysRevD.109.065024] for concrete Unruh-DeWitt-type models, and more recently in [mandrysch2024quantumfieldmeasurementsfewsterverch] for a concrete representation of the FV framework in a Hilbert space.
+Kraus operators in scattering-based frameworks for local measurement in QFT have been derived previously in [46, 39, 37] for concrete Unruh-DeWitt-type models, and more recently in [30] for a concrete representation of the FV framework in a Hilbert space.
 Nevertheless, the operational consequences of the factorisation conditions at the level of Kraus operators have not been investigated.
@@
 IV.2 A concrete model for local field measurements
-We will use a simple model for field measurements (previously considered in [PhysRevD.109.065024]) where a scalar quantum field $\phi$ in Minkowski spacetime is coupled to a free pointer via $P$, with $[Q,P]=i$.
+We will use a simple model for field measurements (previously considered in [37]) where a scalar quantum field $\phi$ in Minkowski spacetime is coupled to a free pointer via $P$, with $[Q,P]=i$.
 We specify the following interaction
@@
 However, to avoid any conceptual problems arising from interpreting the pointer as a relativistic particle—given the well known problems of causality and localisation in relativistic QM (e.g.
-Hegerfeldt’s [hegerfeldt2001particle] and Malament’s [malament1996defense] theorems)—we elect simply to set $H_{\text{pointer}}=0$, viewing the pointer as an abstract, degree of freedom, not tied to spacetime.
+Hegerfeldt’s [24] and Malament’s [29] theorems)—we elect simply to set $H_{\text{pointer}}=0$, viewing the pointer as an abstract, degree of freedom, not tied to spacetime.
 This has no effect on the field Kraus operators, and the change to the $S$-matrix can be found easily.
@@
 $S[f]=\mathcal{T}e^{-i\phi(f)\otimes P}$ (45)
-which we can solve exactly via the Magnus expansion [PhysRevD.109.065024] to give
+which we can solve exactly via the Magnus expansion [37] to give
 $|S[f]=\int\mathrm{d}p\,e^{-ip\phi(f)}e^{-ip^{2}\Delta_{\textsc{r}}(f,f)/2}\otimes\ket{p}\bra{p}$ (46)
@@
 $\Delta_{\textsc{r}}(f,f)=\int\mathrm{d}\mathsf{x}\mathrm{d}\mathsf{y}f(\mathsf{x})\Delta_{\textsc{r}}(\mathsf{x},\mathsf{y})f(\mathsf{y}).$ (47)
-After picking a probe state $\psi(q)$ on the basis of the pointer variable $Q$, using Eq (43) we can compute the Kraus operators for the field [PhysRevD.109.065024]
+After picking a probe state $\psi(q)$ on the basis of the pointer variable $Q$, using Eq (43) we can compute the Kraus operators for the field [37]
 $\Pi_{q}^{\psi}[f]=$ $\frac{1}{\sqrt{2\pi}}\int\mathrm{d}p\,\tilde{\psi}(p)$
@@
 where $P(q)$ is the PVM associated to the possible values of $\phi(f)$ by the spectral theorem.
-In [ozawa1993canonical], Ozawa studies the process of taking such a limit, and shows that a measurement is only repeatable in that limit, in the sense that the outcomes of repeated measurements agree.
+In [36], Ozawa studies the process of taking such a limit, and shows that a measurement is only repeatable in that limit, in the sense that the outcomes of repeated measurements agree.
 Operationally, they consider allowing the interaction to continue for a longer time, say a factor of $T$, and rescaling the pointer units by $1/T$.
@@
 IV.2.1 Regions and sharpness
-To gain some intuition for the bound (50) on how accurately the field values can be probed locally, we can appeal to the results of [PhysRevD.109.065024], where it is shown that for $f$ given by a Gaussian with spacial width $\sigma_{S}$ and time width $\sigma_{T}$,
+To gain some intuition for the bound (50) on how accurately the field values can be probed locally, we can appeal to the results of [37], where it is shown that for $f$ given by a Gaussian with spacial width $\sigma_{S}$ and time width $\sigma_{T}$,
 $\Delta_{\textsc{r}}(f,f)\sim\begin{cases}\frac{1}{\sigma_{T}\sigma_{S}^{3}}&\text{ when }m\sigma_{T}\gg 1\\ \frac{\sigma_{T}}{\sigma_{S}(\sigma^{2}_{S}+\sigma_{T}^{2})}&\text{ when }m=0.\end{cases}$ (52)
@@
 We see that concretely modelling a measurement leads to a channel that would not be easily guessed if avoiding signalling was the only concern.
-Indeed, channels comparable to ours have been guessed, as we will see in the next section, however they generally lack the bound on the sharpness we have found, which comes from the $\Delta_{R}(f,f)$ phase in the $S$-matrix.
+Indeed, channels comparable to ours have been guessed, as we will see in the next section, however they generally lack the bound on the sharpness we have found, which comes from the $\Delta_{\textsc{r}}(f,f)$ phase in the $S$-matrix.
 This phase arose, not from causality constraints, but from explicitly solving for the evolution of the field and pointer.
-As of such, it goes beyond Sorkin’s impossible measurements, as the channel is still causal without the $\Delta_{R}(f,f)$ phase.
-Interestingly, while the appearance of $\Delta_{R}(f,f)$ is not necessary for ruling out signalling of the field channel, it is key in the proof of continuous additivity and thus plays a role in the causality of the dilated channel, see appendix VII.3.
+As of such, it goes beyond Sorkin’s impossible measurements, as the channel is still causal without the $\Delta_{\textsc{r}}(f,f)$ phase.
+Interestingly, while the appearance of $\Delta_{\textsc{r}}(f,f)$ is not necessary for ruling out signalling of the field channel, it is key in the proof of continuous additivity and thus plays a role in the causality of the dilated channel, see Appendix VII.3.
 IV.2.2 The model as a dilation
 Finally, we comment on the connection to previous related work on families of Kraus operators and POVMs for field measurements.
-Our analysis provides a concrete dilation of the $L^{2}$-maps that defines families of Kraus operators in [albertini2023ideal], defined as
+Our analysis provides a concrete dilation of the $L^{2}$-maps that defines families of Kraus operators in [1], defined as
 $K^{f}_{\kappa}(q)=\kappa(q-\phi(f)).$ (53)
@@
 This means that the ‘shape’ of the Kraus operators is specified by the initial state preparation $\psi$ of the probe.
-Similarly, there is a connection to the Kraus operators that are defined in [oeckl2024spectraldecompositionfieldoperators] in the case of instantaneous field measurements, that we will comment on in the next section.
-Lastly, in [mandrysch2024quantumfieldmeasurementsfewsterverch] the authors analyse a linear coupling between two equal mass real scalar field theories, building on [fewster2020quantum, Fewster2023].
+Similarly, there is a connection to the Kraus operators that are defined in [35] in the case of instantaneous field measurements, that we will comment on in the next section.
+Lastly, in [30] the authors analyse a linear coupling between two equal mass real scalar field theories, building on [16, 17].
 By picking a squeezed quasi-local state for the probe field, they show that a Gaussian measurement of the probe leads to an instrument that in the weak coupling limit can be decomposed into a Gaussian measurement and a dephasing channel with certain strength.
@@
 In this section we will examine the relevance of the continuous additivity Eq (31) for the concrete model of field measurements that we introduced in the previous section.
-We will demonstrate the model’s consistency with the framework laid out in section III, and derive the no-signalling and compositional properties of its Kraus operators, generalising those found in [oeckl2024spectraldecompositionfieldoperators].
+We will demonstrate the model’s consistency with the framework laid out in section III, and derive the no-signalling and compositional properties of its Kraus operators, generalising those found in [35].
 These no-signalling identities ensure that this model does not induce impossible operations on the field.
@@
 This provides us with a concrete example of the relationship between no-signalling conditions for dilated and undilated maps.
-Identity (57) was proved in [oeckl2024spectraldecompositionfieldoperators] for an $\epsilon$-family of POVMs, including the spectral decomposition of $\phi(f)$.
+Identity (57) was proved in [35] for an $\epsilon$-family of POVMs, including the spectral decomposition of $\phi(f)$.
 in the case of instantaneous operations defined over a Cauchy surface ($f$ being a purely spacial smearing) and it was proved to be sufficient for enforcing no-signalling condition Eq (21) in this case.
-Note that our results reduce to those in [oeckl2024spectraldecompositionfieldoperators] when we take $\psi$ to be Gaussian with width $\varepsilon$ and $f$ is supported on a spacial hypersurface so that $\Delta_{\textsc{r}}(f,f)=0$.
+Note that our results reduce to those in [35] when we take $\psi$ to be Gaussian with width $\varepsilon$ and $f$ is supported on a spacial hypersurface so that $\Delta_{\textsc{r}}(f,f)=0$.
 In our case, our concrete model describes time-extended operations, so (57) is not going to be sufficient to prove no-signalling and we will need the following additional identity [^17] The connection between this work and a generalised, i.e.
-‘time extended’, version of Oeckl’s results in [oeckl2024spectraldecompositionfieldoperators] using a covariant path-integral description [oeckl2025causalmeasurementquantumfield] is work in progress, see also [oeckl2025localcompositionalmeasurementsquantum]..
+‘time extended’, version of Oeckl’s results in [35] using a covariant path-integral description [34] is work in progress (see also [32])..
 […]
@@
 The fourth line follows from the fact that the argument of $\Phi_{\textsc{a}}$, integrated over $x$, is proportional to the identity i.e.
-using Eq (58) which is shown in the appendix VII.5.
+using Eq (58) which is shown in Appendix VII.5.
 Importantly, this leads to $\Phi_{\textsc{a}}$ dropping out from the expression, and as a result the details of A’s preparation disappear from C’s result.
@@
 It is clear that $f_{t}\rightarrow f\chi_{J_{-}(R)}$ as $t\rightarrow\infty$.
-Finally, we use that the map $f\mapsto\Pi[f]$ is strongly continuous [oeckl2024spectraldecompositionfieldoperators]171717technically, this was proven only for the map between solutions and Kraus operators with $\psi$ Gaussian (see discussion of [oeckl2024spectraldecompositionfieldoperators] in section IV.2)., to see that observables localised to R are updated by the ‘effective’ would-be operators
+Finally, we use that the map $f\mapsto\Pi[f]$ is strongly continuous [35]181818technically, this was proven only for the map between solutions and Kraus operators with $\psi$ Gaussian (see discussion of [35] in section IV.2)., to see that observables localised to R are updated by the ‘effective’ would-be operators
 $\Pi_{q}^{\psi}[f_{\textsc{r}}],\;\;f_{\textsc{r}}(x)=\begin{cases}f&x\in J_{-}(\text{R})\\ 0&x\notin J_{-}(\text{R}),\end{cases}$ (63)
@@
 Importantly, resolving Sorkin’s problem does not rely on this limiting procedure, but on the fact that there is always a choice of hypersurface such that $f_{-}$ will be fully spacelike to A, and so B cannot enable a superuminal signal from A to C (see figure 4).
-After all, even in the case when continuous additivity is not available, impossible operations are excluded by the Hammerstein property, as shown in the appendix VII.1.
+After all, even in the case when continuous additivity is not available, impossible operations are excluded by the Hammerstein property, as shown in Appendix VII.1.
 V.3 Generality of the results
@@
 This is expected as generically only polynomials of first order in the fields lead to no-signalling unitary kicks.
-More details on this will be provided in [simmons_brukner_in_prep].
+More details on this will be provided in [45].
 To summarise, we have shown that for some forms of $S$-matrix, i.e.
@@
 V.4 Measuring observables beyond $\phi(f)$
-We have shown that there exists a very simple solvable model that can measure $\phi(f)$ for any $f$, in a completely causal way, and which recovers the results of [PhysRevD.105.025003, albertini2023ideal, oeckl2024spectraldecompositionfieldoperators], specifically that the Kraus operators should be given by a suitable $L^{2}$-normalised function of $\phi(f)$ (in the spectral calculus sense), which in our case depends on the choice of the state of the probe.
+We have shown that there exists a very simple solvable model that can measure $\phi(f)$ for any $f$, in a completely causal way, and which recovers the results of [28, 1, 35], specifically that the Kraus operators should be given by a suitable $L^{2}$-normalised function of $\phi(f)$ (in the spectral calculus sense), which in our case depends on the choice of the state of the probe.
 However, $\phi(f)$ is a special operator in the algebra: it is unique in that products of smeared fields will generally not generate non-signalling unitary kicks.
-Since the non-signalling of a Gaussian measurement of $\phi(f)$ can be shown to follow from the non-signalling of unitary kicks by $e^{i\phi(f)}$ [PhysRevD.105.025003], it would be surprising if Gaussian (and more general $L^{2}$-normalised) measurements of powers of smeared fields were non-signalling.
-Indeed, it has been shown for a large class of products of smeared fields, that Gaussian measurements are signalling [PhysRevD.105.025003, albertini2023ideal].
+Since the non-signalling of a Gaussian measurement of $\phi(f)$ can be shown to follow from the non-signalling of unitary kicks by $e^{i\phi(f)}$ [28], it would be surprising if Gaussian (and more general $L^{2}$-normalised) measurements of powers of smeared fields were non-signalling.
+Indeed, it has been shown for a large class of products of smeared fields, that Gaussian measurements are signalling [28, 1].
 One way to see this is to note that for sufficiently temporally thin $f$ and in a lattice approximation, $\phi(f)$ can be approximated as
@@
 $+\dots 1\otimes 1\dots\otimes A_{N}.$ (73)
-Compact operators of this form can be measured in a non-signalling way [Borsten:2019cpc] with an update map that projects onto its eigenstate i.e.
+Compact operators of this form can be measured in a non-signalling way [6] with an update map that projects onto its eigenstate i.e.
 a non-destructive measurement.
-However, it is well known from quantum information [PhysRevLett.90.010402, PhysRevA.49.4331] that compact ‘non-local’ operators cannot be measured in a way that does not erase all local information.
+However, it is well known from quantum information [49, 40] that compact ‘non-local’ operators cannot be measured in a way that does not erase all local information.
 Non-local here refers to the fact that they cannot be written in a way that is completely local with respect to the tensor product decomposition e.g.
@@
 […]
-This leads us to conclude that there are two distinct consequences of causality for measurements in QFT: 1) dynamical assumptions preclude the approximations used in the usual von Neumann derivation of ideal measurements, 2) non-local operators must be measured in such a way that the selective state update is destructive [PhysRevA.49.4331, PhysRevLett.90.010402].
+This leads us to conclude that there are two distinct consequences of causality for measurements in QFT: 1) dynamical assumptions preclude the approximations used in the usual von Neumann derivation of ideal measurements, 2) non-local operators must be measured in such a way that the selective state update is destructive [40, 49].
 While a projective measurement or even the measurement model we have studied here has Kraus operators that are simple spectral functions of the observable, and so final state is closely related to the initial state, for non-local observables, we expect the selective version of any causal update rule will lead to a final state that is very different from the original, destroying the initial state.
@@
 Further, we have studied the operational implications for the resulting channels and Kraus operators, leading to causality conditions and the notion of effective Kraus operators.
-This provides a bridge between the abstract channels proposed in [PhysRevD.105.025003, albertini2023ideal, oeckl2024spectraldecompositionfieldoperators] and the rigorous FV framework [fewster2020quantum, PhysRevD.103.025017, fewster_asymptotic_2023].
+This provides a bridge between the abstract channels proposed in [28, 1, 35] and the rigorous FV framework [16, 7, 17].
 The results in this paper hinged on combining quantum information notions of Stinespring and Kraus representations, and abstract notions of local QFT channels, which we introduced section II.
@@
 To reach a sharp measurement, one must replace $f\mapsto f\lambda$, and take $\lambda\rightarrow 0$, amounting to changing the physical coupling (see e.g.
-[mandrysch2024quantumfieldmeasurementsfewsterverch]), and while also taking the probe state to be more sharp in $Q$.
-Similarly, we find that our model exhibits a generalisation of the convolution identity for Kraus operators found in [oeckl2024spectraldecompositionfieldoperators].
+[30]), and while also taking the probe state to be more sharp in $Q$.
+Similarly, we find that our model exhibits a generalisation of the convolution identity for Kraus operators found in [35].
 We’ve argued that this identity is a special case of a more general property of Kraus operators, and follows from the quantum controlled nature of the $S$-matrix in the simple model.
@@
 factorisation of the $S$-matrix across any Cauchy surface.
-This stands in contrast with the conditions proposed in [sorkin1993impossible, Borsten:2019cpc, PhysRevD.105.025003, albertini2023ideal, Fuksa], which are much more global, in the sense that they depend non-trivially on the whole measurement scenario, and do not necessarily reduce to a condition that must hold locally within a given region, independently of the global scenario.
+This stands in contrast with the conditions proposed in [47, 6, 28, 1, 21], which are much more global, in the sense that they depend non-trivially on the whole measurement scenario, and do not necessarily reduce to a condition that must hold locally within a given region, independently of the global scenario.
 Since the stronger form of continuous additivity implies microcausality of the Lagrangian, and microcausality of spacelike Heisenberg fields, we see a restoration of the connection between microcausality and no-signalling, which fits nicely into the AQFT framework.
@@
 Of course, the Stinespring dilation theorem is a general fact about unital CP channels between $C^{*}$-algebras, and so holds in QFT abstractly.
-However, in the literature a causal Stinespring theorem has been eluded to several times [PhysRevD.105.025003, fewster_asymptotic_2023], well summarised by Jubb “It would be useful to construct an explicit dictionary between update maps and specific probe models, and determine if this is possible in general.
+However, in the literature a causal Stinespring theorem has been eluded to several times [28, 17], well summarised by Jubb “It would be useful to construct an explicit dictionary between update maps and specific probe models, and determine if this is possible in general.
 The latter would be analogous to Stinespring’s Dilation theorem, but with the added restriction of locality and causality on the unitary”.
-Relatedly, in this paper we have shown that the abstract maps of [PhysRevD.105.025003, albertini2023ideal, oeckl2024spectraldecompositionfieldoperators], when interpreted as functionals of the function $f$, can be dilated to an $S$-matrix $S[f]$ that obeys factorisation.
+Relatedly, in this paper we have shown that the abstract maps of [28, 1, 35], when interpreted as functionals of the function $f$, can be dilated to an $S$-matrix $S[f]$ that obeys factorisation.
 In this case we found that the Kraus operators have general compositional properties that are directly related to the factorisation of $S[f]$.
@@
 Since we have only focused on a small subset of local $S$-matrices when discussing the properties of the induced Kraus operators, we cannot make general statements.
-However, we believe that this analysis provides hints towards a causal Stinespring theorem, something we will present general results on elsewhere [simmons_brukner_in_prep].
+However, we believe that this analysis provides hints towards a causal Stinespring theorem, something we will present general results on elsewhere [45].
 Acknowledgements.
@@
 We are also grateful to Francisco Calderón and Doreen Fraser for insightful comments on earlier versions of this draft.
+Finally, the authors have benefited from the activities of COST Action CA23115: Relativistic Quantum Information, funded by COST (European Cooperation in Science and Technology) and, in particular, from discussions at EMS-IAMP Spring School: Symmetry and Measurement in Quantum Field Theory at York, United Kingdom.
 VII Appendix
 VII.1 All operations with the Hammerstein property are possible
-In this section we show that operations induced by local $S$-matrices obeying the first three conditions of section III.1 and the Hammerstein property [rejzner2019localitycausalityperturbativeaqft],
+In this section we show that operations induced by local $S$-matrices obeying the first three conditions of section III.1 and the Hammerstein property [42],
 $S[f_{1}+f_{2}+f_{3}]=S[f_{1}+f_{2}]S[f_{2}]^{-1}S[f_{2}+f_{3}],$ (76)
@@
 This implies $\delta=0$, and since we have not made any assumptions about the regions, or $\rho,C$, we can conclude that $S[f]$ is possible if it satisfies the Hammerstein property.
-Specifically, we can see that $\delta=0$ for all conditions implies $S[f]^{\dagger}\mathfrak{A}(\text{C})S[f],\mathfrak{A}(\text{A})]=0$ if $\mathfrak{A}(\text{C}),\mathfrak{A}(\text{A})]=0$, which implies that it is a non-signalling evolution [Borsten:2019cpc].
+Specifically, we can see that $\delta=0$ for all conditions implies $S[f]^{\dagger}\mathfrak{A}(\text{C})S[f],\mathfrak{A}(\text{A})]=0$ if $\mathfrak{A}(\text{C}),\mathfrak{A}(\text{A})]=0$, which implies that it is a non-signalling evolution [6].
 VII.2 Solving linear pointer coupling
@@
 […]
-where $\tilde{\psi}$ is the Fourier transform of $\psi$.
-This simple model satisfies the following convolution property, which was been proven to be relevant for ensuring no-signalling [oeckl2024spectraldecompositionfieldoperators]
+where $\tilde{\psi}$ is the Fourier transform of $\psi$. This simple model satisfies the following convolution property, which was been proven to be relevant for ensuring no-signalling [35]
 $\Pi^{\psi}_{q}[f]=\int\mathrm{d}s\,\Pi^{\psi_{+}}_{q-s}[f_{+}]\Pi^{\psi_{-}}_{s}[f_{-}],$ (93)